### v6

# M. C. Rodriguez

Abstract
We present a preliminar study of the scalar potential of the usual scalars in the Minimal Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model. We will consider the case in which all the usual neutral scalars fields of this model obtain vacuum expectation values. When we allow in our superpotential, only interactions that respect the invariance of the quantum number, ${\cal F}\equiv B+L$, where $B$ is the baryon number while $L$ is the total lepton number. We obtain, as in the minimal $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model without supersymmetry, a Majoron. However, by allowing the violation of this quantum number, we can remove the Majoron in a mechanism similar to that recently presented in the Supersymmetric version of the scheme of Gelmini-Roncadelli. The mass spectrum obtained is in agreement with current experimental data.
Spontaneous breaking of a global symmetry in the Minimal Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model.
M. C. Rodriguez
Grupo de Física Teórica e Matemática Física
Departamento de Física
Universidade Federal Rural do Rio de Janeiro - UFRRJ
BR 465 Km 7, 23890-000, Seropédica - RJ
Brasil
PACS number(s): 12.60. Cn, 12.60. Jv
Keywords: Extension of electroweak gauge sector, Supersymmetric models.
1 Introduction
In the Standard Model (SM) [1, 2], the relationship among the isospin $I_{3}$, the hypercharge $Y$ and the eletromagnetic charge $Q$ is defined by the Gell-Mann-Nishijima equation, which is defined as
$Q=I_{3}+Y.$ (1)
In the SM, we introduce only left-handed neutrinos [^1] We do not to introduce right-handed neutrinos, $\left[\nu_{iR}\sim\left({\bf 1},{\bf 1},0\right)\right]$, in agreement with the two component theory of neutrino theory of Landau [3], Lee and Yang [4] and Salam [5] from 1957, later the helicity of neutrinos was measured experimentally in 1958 and it confirmed this hypothesis [6]. Those right-handed neutrinos are known as sterile neutrinos [7, 8, 9, 10]. as follows [11, 12, 13]
$L_{iL}$ $=$ $\left(\begin{array}[]{c}\nu_{iL}\\ l_{iL}\end{array}\right)\sim\left({\bf 1},{\bf 2},-\frac{1}{2}\right),\,\ l_{iR}\sim\left({\bf 1},{\bf 1},-1\right),\,\ i=1,2,3.$ (4)
We show in parenthesis the transformations properties under the respective factors $(SU(3)_{C},SU(2)_{L},U(1)_{Y})$. The quark sector is defined as
$Q_{iL}$ $=$ $\left(\begin{array}[]{c}u_{iL}\\ d_{iL}\end{array}\right)\sim\left({\bf 3},{\bf 2},+\frac{1}{6}\right),\,\ u_{iR}\sim\left({\bf 3},{\bf 1},+\frac{2}{3}\right),\,\ d_{iR}\sim\left({\bf 3},{\bf 1},-\frac{1}{3}\right).$ (7)
In the SM, we have to introduce the following scalar field $\phi$ in the doublet representation of the $SU(2)_{L}$ [14, 15]
$\phi$ $=$ $\left(\begin{array}[]{c}\phi^{+}\\ \phi^{0}\end{array}\right)\sim\left({\bf 1},{\bf 2},+\frac{1}{2}\right),\,\ \Rightarrow\,\ \tilde{\phi}\equiv\left[\left(i\sigma_{2}\right)\phi^{*}\right]=\left(\begin{array}[]{c}\left(\phi^{0}\right)^{*}\\ -\phi^{-}\end{array}\right)\sim\left({\bf 1},{\bf 2},-\frac{1}{2}\right).$ (13)
This scalar fields get the following vacuum expectation value (VEV)
$\langle\phi\rangle\equiv\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v\end{array}\right),\,\ \langle\tilde{\phi}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}v\\ 0\end{array}\right).$ (19)
The relevant terms in the SM lagrangian are given by
${\cal L}_{SM}={\cal L}^{Y}_{SM}-V(\phi)+{\cal L}^{kin}_{SM}.$ (20)
The charged fermions obtain their masses, via the following Yukawa coupling [11, 12, 13]
${\cal L}^{Y}_{SM}$ $=$ $\left[g^{l}_{ij}\left(\bar{L}_{iL}\phi\right)l_{jR}+g^{d}_{ij}\left(\bar{Q}_{iL}\phi\right)d_{jR}+g^{u}_{ij}\left(\bar{Q}_{iL}\tilde{\phi}\right)u_{jR}+hc\right].$ (21)
The scalar potential of the SM is defined in the following way [14, 15]
$V(\phi)=\mu^{2}|\phi|^{2}+\lambda|\phi|^{4}.$ (22)
Both the $\mu$-parameter and the $\lambda$-parameter are arbitrary and they are free parameters within the SM. For $\mu$-parameter, we can have the following two possibilities [14, 15]
  • $\mu^{2}>0:$ The gauge symmetry $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ is not broken;
  • $\mu^{2}<0:$ We break the $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ gauge symmetry to $SU(3)_{C}\times U(1)_{em}$.
When we impose the condition of extremes
$\left.\frac{\partial V(\phi)}{\partial\phi}\right|_{\langle\phi\rangle=v}=0,$ (23)
the minimum of the potential, in the case where $\mu^{2}<0$, is written in the following way
$v$$\equiv$ $\sqrt{\frac{-\mu^{2}}{\lambda}}=\frac{1}{\sqrt{\sqrt{2}G_{F}}}=246\,\ GeV,$
$G_{F}$ $=$ $\left(1.16639\pm 0.00002\right)\times 10^{-5}\,\ GeV^{-2}\approx\frac{1}{(293\,\ GeV)^{2}}.$ (24)
We make the following shift in the neutral scalar field
$\phi^{0}\equiv\frac{1}{\sqrt{2}}\left(v+h^{0}+iG^{0}\right),$ (25)
under this expansion, the scalar potential of SM become [14]
$V(\phi)$ $=$ $V(\phi^{\prime 0})-\mu^{2}\left(h^{0}\right)^{2}+\lambda vh^{0}\left((h^{0})^{2}+(G^{0})^{2}\right)+\frac{\lambda}{4}\left((h^{0})^{2}+(G^{0})^{2}\right)^{2},$
$V(\phi^{\prime 0})$ $=$ $-\frac{\lambda}{4}v^{4},\,\ |\phi^{\prime 0}|\equiv\frac{v}{\sqrt{2}}=\sqrt{\frac{-\mu^{2}}{2\lambda}}>0.$ (26)
The field $h^{0}$ is CP even field and gets the following mass  [^2] A nice review about this interesting subject can be found in [15].
$M^{2}_{h^{0}}=-2\mu^{2}=2\lambda v^{2}.$ (27)
We must emphasize, the SM do not predict any value for this mass, in fact it can get any values from 00 until a superior limit given by $M_{h^{0}}<1$ TeV [16, 17]. This mechanism is known as Brout-Englert-Higgs mechanism [15, 19].
The mass of the light Higgs boson $h^{0}$ was recently measured experimentally by ATLAS (A Toroidal LHC Apparatus) and CMS (Compact Muon Solenoid) collaborations, at Large Hadron Collider (LHC). Their measurements for the mass of the Higgs is [20, 21]
$M_{h^{0}}=\left(125.20\pm 0.11\right)\,\ GeV.$ (28)
After breaking the gauge symmetry the gauge boson get mass, through the following term
${\cal L}^{kin}_{SM}$ $=$ $\left(D_{m}\phi\right)^{\dagger}\left(D^{m}\phi\right),$
$D_{m}\phi$ $=$ $\left[\partial_{m}-ig\sum_{i=1}^{3}\left(\frac{\sigma^{i}}{2}\right)W^{i}_{m}-ig^{\prime}\left(\frac{1}{2}\right)W_{m}\right]\phi,$ (29)
where $W^{i}_{m}$ is the $SU(2)_{L}$ gauge bosons and $W_{m}$ is the $U(1)_{Y}$ gauge boson, while $g$ and $g^{\prime}$ are the gauge constant of the group $SU(2)_{L}$ and $U(1)_{Y}$, respectively. We obtain in addition to photon $A_{m}$, two massive bosons $W^{\pm}_{m}$ and $Z^{0}_{m}$ and their masses have the following expressions [^3] The numerical value of the first relation below was obtained from [22] [11, 12]
$M^{2}_{W^{\pm}}$ $=$ $\frac{g^{2}v^{2}}{4}=\left[\left(80.3505\pm 0.0077\right)\,\ GeV\right]^{2},$
$M^{2}_{Z^{0}}$ $=$ $\left(\frac{g^{2}+(g^{\prime})^{2}}{4}\right)v^{2}=\frac{g^{2}v^{2}}{4}\left(1+\tan^{2}\theta_{W}\right)=\frac{M^{2}_{W^{\pm}}}{\cos^{2}\theta_{W}},$
$W^{\pm}_{m}$ $=$ $\frac{1}{\sqrt{2}}\left(W^{1}_{m}\mp W^{1}_{m}\right),\,\ \left(\begin{array}[]{c}Z^{0}_{m}\\ A_{m}\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta_{W}&-\sin\theta_{W}\\ \sin\theta_{W}&\cos\theta_{W}\end{array}\right)\left(\begin{array}[]{c}W_{3m}\\ W_{m}\end{array}\right),$ (36)
where $\theta_{W}$ is the Weinberg angle and it is defined as follows
$\tan\theta_{W}\equiv\frac{g^{\prime}}{g},\,\ \Rightarrow\,\ e=g\sin\theta_{W}=g^{\prime}\cos\theta_{W}.$ (38)
The SM Lagrangian have four free parameters: $g,g^{\prime},\mu^{2}$ and $\lambda$ in such a way, we can reproduce the old low-energy data [11, 12, 14].
The average of the experimental values for the gauge bosons masses are [20, 22]
$\left(M_{W^{\pm}}\right)_{ave}$ $=$ $\left(80.4133\pm 0.0080\right)GeV,$
$\left(M_{Z^{0}}\right)_{ave}$ $=$ $\left(91.1875\pm 0.0021\right)GeV,$ (39)
while for the Weinberg angle, $\theta_{W}$, we have
$\sin^{2}\theta_{W}$ $=$ $0.2324\pm 0.0012;\,\ \Rightarrow\,\ \cos^{2}\theta_{W}\equiv\left(\frac{M_{W^{\pm}}}{M_{Z^{0}}}\right)^{2}\sim 0.77$ (40)
Today we know neutrinos are massive particles [23, 24, 25, 26, 27, 28]. If we introduce sterile neutrinos in SM, $\nu_{iR}$, we can write the following Dirac mass term for neutrinos
${\cal L}^{\nu}_{\phi,L}=y^{\nu}_{ij}\left[\left(\bar{L}_{iL}\tilde{\phi}\right)\nu_{jR}+hc\right],$ (41)
where $\tilde{\phi}$ is defined in our Eq.(25). This mass term conserve lepton number defined in our Eq.(43) and in this case the neutrinos are Dirac-type particles. A priori, to solve this problem, we can also to introduce the following Majorana mass term for the left-handed neutrinos [25]
$m^{M}_{ij}\left(\overline{\nu^{c}_{iL}}\nu_{jL}+hc\right).$ (42)
This term break lepton number conservation, remember the SM respects the following accidental global symmetry [8]
$G^{global}_{SM}\equiv U(1)_{B}\times U(1)_{L_{e}}\times U(1)_{L_{\mu}}\times U(1)_{L_{\tau}},$ (43)
where $B$ is the baryon number and $L_{e},L_{\mu}$ and $L_{\tau}$ are the lepton number of each lepton.
An interesting way to generate the Eq.(42), is to add the following term for the Yukawa coupling, defined by Eq.(21), lagrangian of SM [13]
$\overline{L^{c}_{iL}}L_{jL}\sim\left({\bf 1},{\bf 1\oplus 3},-1\right);\,\ L^{c}_{iL}=C\bar{\psi}^{T}\sim\gamma_{2}f^{*}=\left(\begin{array}[]{c}\nu^{c}_{iL}\\ l^{c}_{iL}\end{array}\right)\sim\left({\bf 1},{\bf 2},+\frac{1}{2}\right).$ (44)
for $f=\nu$ and $e$. Therefore, to provide mass to neutrinos, we need to introduce new fields within the SM context. We can add to the SM singlets and also the triplets, of scalars and/or fermions, in representation of $SU(2)_{L}$. When we introduce
  • Extra fermions in singlet, we generate masses for neutrinos by the known type I Seesaw mechanism [29, 30, 31];
  • Extra scalars in triplets, it arise type II Seesaw mechanism [32, 33, 34, 35, 36];
  • Extra fermions in triplets, we obtain type III Seesaw mechanism [37].
This new fields, will generate Leptogenesis [38, 39, 40] an attractive scenario to explain the baryon asymmetry of the Universe [41, 42].
We can break the lepton number symmetry [^4] The total lepton number $L$ is defined in terms of $L_{e},L_{\mu}$ and $L_{\tau}$, see Eq.(43), in the following way $L=L_{e}+L_{\mu}+L_{\tau}$. when we introduce the new scalar $\Delta$ in the triplet representation [^5] The triplet $\Delta$ carries the lepton number of $2$., see Eq.(44), of the group $SU(2)_{L}$
$\Delta=\left(\begin{array}[]{cc}\frac{h^{+}}{\sqrt{2}}&H^{++}\\ \Delta^{0}&-\frac{h^{+}}{\sqrt{2}}\end{array}\right)\sim\left({\bf 1},{\bf 3},+1\right),\,\ \langle\Delta\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&0\\ V_{\Delta}&0\end{array}\right).$ (45)
This scalar allows us to write the following Yukawa interactions for neutrinos [13, 43]
${\cal L}^{\nu}_{\Delta,L}$ $=$ $g^{\nu}_{ij}\left[\left(\overline{L^{c}_{iL}}\Delta L_{jL}\right)+hc\right]$
$=$ $g^{\nu}_{ij}\left[\overline{\nu^{c}_{iL}}\nu_{jL}\Delta^{0}+\left(\overline{\nu^{c}_{iL}}l_{jL}-\overline{l^{c}_{iL}}\nu_{jL}\right)\frac{h^{+}}{\sqrt{2}}+\overline{l^{c}_{iL}}l_{jL}H^{++}+hc\right],$
the Yukawa coupling $h_{ij}$ is the $3\times 3$ complex symmetric matrix. Those scalars $h^{+},\,\ H^{++}$ and $\Delta^{0}$ are known as bileptons [44]. The first term in the above equation is exactly our Eq.(42), where we can identify
$m^{M}_{ij}=g^{\nu}_{ij}\frac{V_{\Delta}}{\sqrt{2}}.$ (47)
It is know as Gelmini-Roncadelli scheme [43], it is type II Seesaw mechanism [32, 33, 34, 35, 36] and neutrinos are Majorana-type particles rather than being a Dirac-type particle such as all charged fermions. The double beta decay experiments, is defined in the following way
$(A,Z)\rightarrow(A,Z+2)+2e^{-},$ (48)
can show if the neutrinos are Majorana particles [45] or Dirac particles [46], see our Eq.(41), more information on this subject can be seen at [9, 10, 25, 47, 48, 49].
A useful parameter for placing strong limits in physics beyond the SM is the so-called $\rho$-parameter, this parameter is defined as follows
$\rho\equiv\left(\frac{M^{2}_{W^{\pm}}}{M^{2}_{Z^{0}}\cos^{2}\theta_{W}}\right),\,\ \Rightarrow\rho^{SM}=1,$ (49)
which is in perfect agreement with its experimental value given by [20]
$\rho_{\rm EXP}=0.9998\pm 0.0008.$ (50)
Using the new scalar triplet, $\Delta$ defined in our Eq.(45), we can in addition to generating masses for neutrinos, jointly explain the recent measurement for $W$-boson mass, obtained by CDF collaboration [50]
$\left(M_{W^{\pm}}\right)_{CDF}=\left(80.4335\pm 0.0094\right)\,\ GeV,$ (51)
because the masses of the gauge bosons in the scheme of Gelmini-Roncadelli are written as follows [51, 52, 53]
$M^{2}_{W^{\pm}}$ $=$ $\frac{g^{2}}{4}\left(v^{2}+2V^{2}_{\Delta}\right)=\frac{g^{2}v^{2}}{4}\left(1+2R\right),\,\ R=\left(\frac{V_{\Delta}}{v}\right)^{2},$
$M^{2}_{Z^{0}}$ $=$ $\left(\frac{g^{2}+(g^{\prime})^{2}}{4}\right)\left(v^{2}+4V^{2}_{\Delta}\right)=\frac{g^{2}v^{2}}{4\cos^{2}\theta_{W}}\left(1+4R\right).$ (52)
We can explain the new measurement for the $W$-boson mass if $V_{\Delta}\approx 10$ GeV, as can be easily seen in our Fig.(1).
Figure 1: We show the prediction about the behaviour for the gauge boson mass, $M_{W^{\pm}}\equiv M_{W}$, in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $V_{\Delta}$ (black curve) and we take $v=246$ GeV. The straight lines in red, green and blue are the values show in our Eqs.(LABEL:wsmval,39,51), respectively.
On the other hand, the $\rho$-parameter in the scheme of Gelmini-Roncadelli, see Eq.(47), becomes [54]
$\rho$ $=$ $\frac{1+2R}{1+4R}\approx 1-2R.$ (53)
With this expression, it is clear that we can explain the current experimental value of the $\rho$-parameter when $V_{\Delta}\approx 4.92$ GeV, see our Fig.(2).
Figure 2: We show the prediction about the behaviour for the $\rho$-parameter in the scheme of Gelmini-Roncadelli, see our Eq.(53), as a function of the Triplet VEV $V_{\Delta}$ (black curve) and we take $v=246$ GeV. The straight lines in red, green and blue are the following values $\rho^{inf}=\rho^{cen}-0.0008$, $\rho^{cen}=0.9998$ and $\rho^{SM}=1$ (as defined in our Eq.(49)), respectively.
In the scheme of Gelmini-Roncadelli, the scalar potential is written in the following form [55, 56]
$V_{G}(\phi,\Delta)$ $=$ $\mu^{2}(\phi^{\dagger}\phi)+M^{2}Tr(\Delta^{\dagger}\Delta)+\lambda_{1}(\phi^{\dagger}\phi)^{2}+\lambda_{2}\left[Tr(\Delta^{\dagger}\Delta)\right]^{2}$ (54)
$+$ $\lambda_{3}Tr[(\Delta^{\dagger}\Delta)(\Delta^{\dagger}\Delta)]+\lambda_{4}(\phi^{\dagger}\phi)Tr(\Delta^{\dagger}\Delta)$
$+$ $i\lambda_{5}\epsilon_{ijk}(\phi^{\dagger}\sigma^{i}\phi)(\Delta^{j\dagger}\Delta^{k}).$
Now, if we make the shift defined in our Eq.(25) together with the following expansion
$\Delta^{0}=\frac{1}{\sqrt{2}}\left(V_{\Delta}+R_{\Delta}+iI_{\Delta}\right).$ (55)
We get the following constraint equations
$\frac{t_{H}}{v}$ $=$ $\left(\mu^{2}+\lambda_{1}v^{2}+\frac{\lambda_{3}}{2}v^{2}_{\Delta}-\frac{\lambda_{4}}{2}v^{2}_{\Delta}\right)=0,$
$\frac{t_{\Delta}}{V_{\Delta}}$ $=$ $\left(M^{2}+\lambda_{2}v^{2}_{\Delta}+\frac{\lambda_{3}}{2}v^{2}_{\Delta}-\frac{\lambda_{4}}{2}v^{2}_{\Delta}+\lambda_{5}v^{2}_{\Delta}\right)=0.$ (56)
After carrying out the shift, defined in our Eqs.(25,55), we obtain the following mass matrices for the neutral scalars in the scheme of Gelmini-Roncadelli
$\frac{1}{2}\left(\begin{array}[]{cc}R_{H}&R_{\Delta}\end{array}\right)M^{2}_{R}\left(\begin{array}[]{c}R_{H}\\ R_{\Delta}\end{array}\right)+\frac{1}{2}\left(\begin{array}[]{cc}I_{H}&I_{\Delta}\end{array}\right)M^{2}_{I}\left(\begin{array}[]{c}I_{H}\\ I_{\Delta}\end{array}\right).$ (63)
For the scalar of CP even, we get the following mass matrix
$M^{2}_{R}=\left(\begin{array}[]{cc}2\lambda_{1}v^{2}&(\lambda_{3}-\lambda_{4})vV_{\Delta}\\ (\lambda_{3}-\lambda_{4})vV_{\Delta}&2\lambda_{2}V^{2}_{\Delta}\end{array}\right).$ (64)
Where we get the following analytical results $det\left(M^{2}_{R}\right)\neq 0$ and $Tr\left(M^{2}_{R}\right)\neq 0$, therefore we have two massive scalars. The lightest Higgs, we will represent as $h^{0}$, and we also have another heavier Higgs, which we will denote as $H^{0}$. The mass of the lightest scalar, is the same as the Higgs boson of SM, compare with Eq.(27), is [^6] The complete expression for the mass of these scalars can be found in [57]. The mass of heavier Higgs, $H^{0}$, we will present below.
$M^{2}_{h^{0}}\approx 2\lambda_{1}v^{2}.$ (65)
The mass $M_{h^{0}}$ as function of the $\lambda_{1}$-parameter is shown in our Fig.(3). Where, we can see that we can get the central data of the lightest Higgs mass if we assume $\lambda_{1}\approx 0.1295$.
Figure 3: We show the prediction about the behaviour for the lightest Higgs mass, $M_{h^{0}}$, in the scheme of Gelmini-Roncadelli, see our Eq.(65), as function of the $\lambda_{1}$-parameter (black curve) and $v=246$ GeV. The $M^{exp}_{h^{0}}$ is the central experimental values (red line), given at our Eq.(28), and $M^{sup}_{h^{0}}=M^{exp}_{h^{0}}+0.11$ (green line) while $M^{inf}_{h^{0}}=M^{exp}_{h^{0}}-0.11$ (blue line).
For the case of CP odd sector, we obtain the following mass matrix
$M^{2}_{I}=\left(\begin{array}[]{cc}\frac{t_{H}}{v}&0\\ 0&\frac{t_{\Delta}}{V_{\Delta}}\end{array}\right).$ (66)
We get two Goldstone bosons, because $det\left(M^{2}_{I}\right)=0$ and $Tr\left(M^{2}_{I}\right)=0$, one of them is the same $G^{0}$ obtained in SM. The extra Goldstone boson is the Majoron, $M^{0}\equiv I_{\Delta}$ [43]. The Majoron, $M^{0}$, can be detected experimentally through the following process [47, 48, 49]
$(A,Z)\rightarrow(A,Z+2)+2e^{-}+M^{0}.$ (67)
The existence of a non-massive Majoron, $M^{0}$, is excluded by experimental data [^7] Decay $Z^{0}\rightarrow M^{0}+X$. [58]. One possible solution for this problem, is to make the Majoron, $M^{0}$, invisible to the leptons and quarks, but we will not use this solution in this article.
There are another three possible solutions to remove the non-massive Majoron [56]
  • By Higgs mechanism, $R_{\Delta}$ and $I_{\Delta}$ have to be absorbed;
  • Giving a large mass to the Majoron ($M^{0}$) without VEV;
  • Giving a large mass to the Majoron ($M^{0}$) with tiny VEV.
Let’s quickly review the type of the last solution, where we need to add the following new term in the scalar potential  [^8] We will denote it as scheme of Ma. [55, 56]
$V(\phi,\Delta)$ $=$ $V_{G}(\phi,\Delta)+\left(\mu_{M}\left(\phi^{\dagger}\Delta\tilde{\phi}\right)+hc\right).$ (68)
Where $V_{G}(\phi,\Delta)$ is defined in our Eq.(54) and $\tilde{\phi}$ is defined in our Eq.(25). $\langle\Delta\rangle$ give mass only for neutrinos, review our Eq.(47), then it has tiny VEV, which is precisely the third case mentioned above. In the scheme of Ma, our Eq.(69) become
$M^{2}_{R}=\left(\begin{array}[]{cc}2\lambda_{1}v^{2}&(\lambda_{3}-\lambda_{4})vV_{\Delta}+2\sqrt{2}\mu_{M}v\\ (\lambda_{3}-\lambda_{4})vV_{\Delta}+2\sqrt{2}\mu_{M}v&2\lambda_{2}V^{2}_{\Delta}-\frac{\mu_{M}}{\sqrt{2}}\frac{v^{2}}{V_{\Delta}}\end{array}\right).$ (69)
The mass spectrum, we get beyond the lightest Higgs, $h^{0}$, given in Eq.(65), we obtain the heavier Higgs field, $H^{0}$, and its mass is given by
$M^{2}_{H^{0}}\approx\frac{\mu_{M}}{\sqrt{2}}\frac{v^{2}}{V_{\Delta}}.$ (70)
$M_{H^{0}}$ as function of, the VEV, $V_{\Delta}$ is shown in our Fig.(4) and we obtain masses within experimental limits for the heavy Higgs [59]
$400\,\ GeV\,\ \leq M_{H^{0}}\leq 1000\,\ GeV,$ (71)
if $V_{\Delta}\leq 3$ GeV and $\mu\geq 10^{0}$ GeV.
Figure 4: We show the prediction about the behaviour for the heavier Higgs, $M_{H^{0}}$, in the scheme of Ma, see our Eq.(70), as a function of the Triplet VEV $V_{\Delta}$, and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ show in the box in several colors. $M^{exp}_{H^{0}}$ is the lower experimental values given at our Eq.(71) in orange line.
The mass matrix for neutral CP odd, we get one massive pseudoscalar, $A^{0}$, [55, 56]
$M^{2}_{A^{0}}\approx\frac{\mu_{M}}{\sqrt{2}}\left(\frac{4v^{2}_{\Delta}+v^{2}}{V_{\Delta}}\right)=M^{2}_{H^{0}}+2\sqrt{2}\mu_{M}V_{\Delta},$ (72)
therefore $M_{A^{0}}>M_{H^{0}}$. We show $M_{A^{0}}$ as function of, the VEV, $V_{\Delta}$ in our Fig.(5) and again our results for the mass of $M_{A^{0}}$ is in agreement experimental limits for the pseudoscalar, $A^{0}$, given by the following constraint [59]
$400\,\ GeV\,\ \leq M_{A^{0}}\leq 1000\,\ GeV,$ (73)
if $V_{\Delta}\leq 3$ GeV and $\mu\geq 10^{0}$ GeV.
Figure 5: We show the prediction about the behaviour for the pseudoscalar, $M_{A^{0}}$, in the scheme of Ma, see our Eq.(72), as a function of the Triplet VEV $V_{\Delta}$, and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ show in the box in several colors. $M^{exp}_{A^{0}}$ is the lower experimental values given at our Eq.(73) in orange line.
In the scheme of Ma, see our Eq.(68), we also have charged scalars and their masses are given by the following relations [60, 79]
$M^{2}_{h^{\pm}}=M^{2}_{H^{0}}-\frac{\lambda_{5}}{4}v^{2},\,\ M^{2}_{h^{\pm\pm}}=M^{2}_{H^{0}}-\frac{\lambda_{5}}{2}v^{2}.$ (74)
We have the following hierarchy if $\lambda_{5}>0$
[…]
while for the case of $\lambda_{5}<0$ we have
[…]
Our numerical results satisfy the following experimental constraints $M_{h^{\pm}}>155$ GeV [20] and $M_{h^{\pm\pm}}>445$ GeV [62]. We want to emphasize that the masses of the charged scalars for the case $\lambda_{5}<0$ are heavier than in the opposite case.
Figure 6: We show the prediction about the behaviour for the masses, $M_{H^{0}}$ (blue curve), and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ show in the box in several colors. The $M_{h^{\pm}}$ (green curve) and $M_{h^{\pm\pm}}$ (red curve) when $\lambda_{5}>0$.
Figure 7: We show the prediction about the behaviour for the masses $M_{h^{\pm}}$ (green curve) and $M_{h^{\pm\pm}}$ (red curve) when $\lambda_{5}<0$ and $M_{H^{0}}=654.150$ GeV. The values contained in this graph are in agreement with the analytical results obtained by Eqs.(75,LABEL:numvall5n).
Supersymmetry, it is more known as SUSY, arose in theoretical papers more than 30 years ago independently by Golfand and Likhtman [63], Volkov and Akulov [64] and Wess and Zumino [65]. The supersymmetry algebra was introduced in [63]  [^9] They also constructed the first four-dimensional field theory with supersymmetry, (massive) quantum electrodynamics (SQED) of spinors and scalars. We will call it as Golfand-Likhtman model as suggested by Shiffman in his article [66].. The Volkov-Akulov article, however, deals only with fermions and its name is “Is the Neutrino a Goldstone Particle? [^10] It is the base of supergravity.. For interested people interested in the history of supersymmetry see [67, 68] and for the earliest Supersymmetric Standard Model see [69, 70], while anyone interested, how to work with supersymmetry see the great review article [71].
The Supersymmetric Standard Model (MSSM) is a good candidate to be the physics beyond the SM. In this model we can solve the hierarchy problem as well to explain the Higgs Masses [^11] The current status of the search for supersymmetry is presented in reference [72]. [73, 74]. In the MSSM, as in the SM, neutrinos are massless. If we want to give mass to the neutrinos, we must break $R$-Parity. In this mechanism, the masses generated are a combination of type-I [29, 30, 31] and type-III Seesaw [37] mechanism. Another interesting result that arises when we break $R$-Parity is to explain the new $W$ mass get from the CDF collaboration [75].
There is also, the Minimal $R$-Symmetric extension of the Minimal Superymmetric Standard Model, known as MRSSM [76, 80, 78]. The field content of the MRSSM is enlarge compared to the MSSM as we shown in our Tab.(1)and the triplet $\hat{T}$ is defined as [80, 78]
$\hat{T}$ $=$ $\left(\begin{array}[]{cc}\frac{\hat{T^{0}}}{\sqrt{2}}&\hat{T}^{+}\\ \hat{T}^{-}&\frac{\hat{T^{0}}}{\sqrt{2}}\end{array}\right).$ (79)
SuperfieldField $(SU(3)_{C},SU(2)_{L})_{U(1)_{Y}}$ $U(1)_{R}$
$\hat{Q}_{L}$ $({\bf 3},{\bf 2})_{1/6}$ 1
$\hat{U}_{R}$ $({\bf\bar{3}},{\bf 1})_{-2/3}$ 1
$\hat{D}_{R}$ $({\bf\bar{3}},{\bf 1})_{1/3}$ 1
$\hat{L}_{L}$ $({\bf 1},{\bf 2})_{-1/2}$ 1
$\hat{E}_{R}$ $({\bf 1},{\bf 1})_{1}$ 1
$\hat{S}$ $({\bf 1},{\bf 1})_{0}$ 0
$\hat{T}$ $({\bf 1},{\bf 3})_{0}$ 0
$\hat{{\cal O}}$ $({\bf 8},{\bf 1})_{0}$ 0
$\hat{H}_{1}$ $({\bf 1},{\bf 2})_{1/2}$ 0
$\hat{H}_{2}$ $({\bf 1},{\bf 2})_{-1/2}$ 0
$\hat{T}$ $({\bf 1},{\bf 3})_{0}$ 0
$\hat{S}$ $({\bf 1},{\bf 1})_{0}$ 0
$\hat{R}_{1}$ $({\bf 1},{\bf 2})_{-1/2}$ 2
$\hat{R}_{2}$ $({\bf 1},{\bf 2})_{+1/2}$ 2
Table 1: Matter and $R$-charges in the $R$-symmetric supersymmetric model, MRSSM.
The superpotential of this model is [79]
$W=\left(\mu_{1}+\lambda_{1}\hat{S}\right)\hat{H}_{1}\hat{R}_{1}+\left(\mu_{2}+\lambda_{2}\hat{S}\right)\hat{H}_{2}\hat{R}_{2}+\Lambda_{1}\hat{R}_{1}\hat{T}\hat{H}_{1}+\Lambda_{2}\hat{R}_{2}\hat{T}\hat{H}_{2}.$ (80)
We can generate Majorana mass for neutrinos through the operator
$H_{2}H_{2}L_{iL}L_{jL}.$ (81)
The mass of $W$ is given by [80, 78]
$M^{2}_{W}=\frac{g^{2}}{4}\left(v^{2}+4v^{2}_{T}\right)=\frac{g^{2}v^{2}}{4}\left(1+R\right),\,\ R=\left(\frac{2v_{T}}{v}\right)^{2},$ (82)
We can explain the new measurement for the $W$-boson mass, given at Eq.(51), if $|v_{T}|\leq 4$ GeV [80, 78] and see our Fig.(8). The mass expression for $Z$ gauge boson has the same expression as presented in our Eq.(LABEL:wsmval). The $\rho$-parameter is [80, 78]
$\rho=\left(1+R\right),\,\ R=\left(\frac{2v_{T}}{v}\right)^{2},$ (83)
and see our Eq.(9). The mass spectrum in Higgs sector was presented [80, 79, 80, 81].
Figure 8: We show the prediction about the behaviour for the gauge boson mass $W^{\pm}$ in MRSSM, see our Eq.(82), as a function of the Triplet VEV $v_{T}$. The straight lines in blue, red and green are the values show in our Eqs.(LABEL:wsmval,39,51), respectively.
Figure 9: We show the prediction about the behaviour for the $\rho$-parameter in MRSSM, see our Eq.(83), as a function of the Triplet VEV $v_{T}$. Where we use the following values $\rho^{cen}=0.9998$, $\rho^{lim}_{sup}=\rho^{cen}+0.0008$ and $\rho^{SM}=1$ as defined in our Eq.(49).
There are also the supersymmetric version of the scheme of Gelmini-Roncadelli  [^12] We will call this model as SUSYGR by short. [82, 83, 84, 85, 86], where as in the MSSM, we have two Higgs in the doublet representations  [^13] We need also to introduce their superpartners, known as higgsinos. They are put in the chiral superfields [71, 73, 74].
$H_{1}$ $=$ $\left(\begin{array}[]{c}h^{0}_{1}\\ h^{-}_{1}\end{array}\right)\sim\left({\bf 1},{\bf 2},-\frac{1}{2}\right),\,\ \langle H_{1}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}v_{1}\\ 0\end{array}\right),$ (88)
$H_{2}$ $=$ $\left(\begin{array}[]{c}h^{+}_{2}\\ h^{0}_{2}\end{array}\right)\sim\left({\bf 1},{\bf 2},+\frac{1}{2}\right),\,\ \langle H_{2}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v_{2}\end{array}\right).$ (93)
We also add one scalar in the triplet representation ${\bf T}_{1}=(T_{1},T_{2},T_{3})$ and another one in the anti-triplet representation ${\bf T}_{2}=(\bar{T}_{1},\bar{T}_{2},\bar{T}_{3})$, where
$\Delta_{1}$ $=$ $\left(\begin{array}[]{cc}\Delta^{0}_{1}&-\frac{h^{+}_{1}}{\sqrt{2}}\\ -\frac{h^{+}_{1}}{\sqrt{2}}&-H^{++}_{1}\end{array}\right)\sim\left({\bf 1},{\bf 3},+1\right),\,\ \langle\Delta_{1}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}V_{\Delta_{1}}&0\\ 0&0\end{array}\right),$ (98)
$\Delta_{2}$ $=$ $\left(\begin{array}[]{cc}\Delta^{--}_{2}&-\frac{h^{-}_{2}}{\sqrt{2}}\\ -\frac{h^{-}_{2}}{\sqrt{2}}&-\Delta^{0}_{2}\end{array}\right)\sim\left({\bf 1},{\bf\bar{3}},-1\right),\,\ \langle\Delta_{2}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&0\\ 0&V_{\Delta_{2}}\end{array}\right).$ (103)
The superpotential of the SUSYGR is written as follows [82, 83, 84, 85, 86]
$W^{GR}_{SUSY}$ $=$ $W^{MSSM}_{RC}+\mu_{\Delta}\left(\hat{\Delta}_{1}\hat{\Delta}_{2}\right)+f^{N}_{ij}Tr\left[\hat{L}_{i}\hat{\Delta}_{1}\hat{L}_{j}\right]+\lambda_{1}Tr\left[\hat{H}_{1}\hat{\Delta}_{1}\hat{H}_{1}\right]$
$+$ $\lambda_{2}Tr\left[\hat{H}_{2}\hat{\Delta}_{2}\hat{H}_{2}\right],$
$W^{MSSM}_{RC}$ $=$ $\mu\left(\hat{H}_{1}\hat{H}_{2}\right)+f^{l}_{ij}\left(\hat{H}_{1}\hat{L}_{i}\right)\hat{E}_{j}+f^{d}_{ij}\left(\hat{H}_{1}\hat{Q}_{i}\right)\hat{D}_{j}+f^{u}_{ij}\left(\hat{H}_{2}\hat{Q}_{i}\right)\hat{U}_{j}.$
Where $\left(\hat{\Delta}_{1}\hat{\Delta}_{2}\right)\equiv\epsilon_{\alpha\beta}\hat{\Delta}_{1}^{\alpha}\hat{\Delta}_{2}^{\beta}$ and $W^{MSSM}_{RC}$ is the term already well known from MSSM with $R$ parity conservation and you can see this term in [74]. Therefore the mass mechanism for the neutrinos is the same as in the scheme of Gelmini-Roncadelli and it is presented in [82, 85].
The scalar field $\Delta_{2}$ is not necessary to give mass for any fermion in the SUSYGR. The terms proportional to $\lambda_{1}$ and $\lambda_{2}$ explicitly break lepton number conservation if we assign $L=-2\,\ (2)$ to the triplet $\Delta_{1}\,\ (\Delta_{2})$ [82]. Therefore, these two parameters break the lepton number conservation, $L$, and they will allow us to implement the Scheme of Ma, see our Eq.(68), in the SUSYGR.
The masses of the gauge bosons in the SUSYGR can be written in the following way
$M^{2}_{W^{\pm}}$ $=$ $\frac{g^{2}}{4}\left[v^{2}+2\left(V^{2}_{\Delta_{1}}+V^{2}_{\Delta_{2}}\right)\right]=\frac{g^{2}v^{2}}{4}\left[1+2\left(R_{1}+R_{2}\right)\right],$
$M^{2}_{Z^{0}}$ $=$ $\frac{g^{2}}{4\cos^{2}\theta_{W}}\left[v^{2}+4\left(V^{2}_{\Delta_{1}}+V^{2}_{\Delta_{2}}\right)\right]=\frac{g^{2}v^{2}}{4\cos^{2}\theta_{W}}\left[1+4\left(R_{1}+R_{2}\right)\right],$
$\rho$ $=$ $\frac{1+2\left(R_{1}+R_{2}\right)}{1+4\left(R_{1}+R_{2}\right)}\approx 1-2\left(R_{1}+R_{2}\right),$
$R_{1}$ $=$ $\left(\frac{V_{\Delta_{1}}}{v}\right)^{2},\,\ R_{2}=\left(\frac{V_{\Delta_{2}}}{v}\right)^{2}.$ (106)
Our numerical results are show in Figs.(10,11). The usual scalar sector of this model is presented in [57, 60, 79, 83]. The resolution of these two problems at the same time in this model, we believe, should be better studied later in this SUSYGR.
Figure 10: We show the prediction about the behaviour for the gauge boson mass $W^{\pm}$ in SUSYGR, see our Eq.(106), as a function of the Triplet VEV $u_{2}\equiv V_{\Delta_{2}}$, and some values of $u_{1}\equiv V_{\Delta_{1}}$ show in the box in several colors and $v=246$ GeV. The straight lines in brown, orange and cyan are the values show in our Eqs.(LABEL:wsmval,39,51), respectively.
Figure 11: We show the prediction about the behaviour for the $\rho$-parameter in SUSYGR, see our Eq.(106), as a function of the Triplet VEV $u_{2}\equiv V_{\Delta_{2}}$, and some values of $u_{1}\equiv V_{\Delta_{1}}$ show in the box in several colors and $v=246$ GeV. Where we use the following values $\rho^{cen}=0.9998$, $\rho^{inf}=\rho^{cen}-0.0008$ and $\rho^{SM}=1$ as defined in our Eq.(49).
Models with the following gauge symmetry
$SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ (107)
are known as $331$ for short. They are interesting possibilities for the physics at the TeV scale due the following nice aspects [87, 88, 89]
  1. 1.
    The family number must be a multiple of three in order to cancel anomalies [88];
  2. 2.
    Why $\sin^{2}\theta_{W}<\frac{1}{4}$ is observed [88];
  3. 3.
    It is the simplest model that includes bileptons of both types: scalars and vectors ones, see our Eq.(LABEL:lq);
  4. 4.
    The models have a scalar sector similar to the two Higgs doublets Model, see Sec.(2) or our Eqs.(114,120,126,132), which allow to predict the quantization of the electric charge [90];
  5. 5.
    It solves the strong $CP$-problem [91];
  6. 6.
    The model has several sources of CP violation [92, 93, 94].
We will concentrate in the model presented in Ref. [88], and the solution to the anomaly for the mass of the $W$-boson in the $331$ was recently presented in the following reference [95]. The scalar sector of the minimal $331$ model, $m331$, was presented in [51, 88, 96, 97]. We show the most general scalar potential in our Eq.(200). For example the terms proportional to $(\eta S\eta^{\dagger}\eta)$ and $(\chi\rho SS)$ does not conserve lepton number conservation, and they are very similar to $\mu$, implemented the scheme of Ma, presented in our Eq.(68).
A few years ago, it was made a preliminary study about the Majoron in the $m331$, in this article they do not used interactions that break lepton number conservation [51]. Therefore, they do not remove the Majoron, $M^{0}$. However in the $m331$ the Majoron, $M^{0}$, does not couple, at the tree level, neither to charged leptons nor to the quarks, then we avoid the so dangerous decay $Z^{0}\rightarrow R_{\Delta}+M^{0}$ [51, 98].
The minimal supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model, or it is more known as MSUSY331 for shortness, was built in [99, 100, 101, 102, 103]. An interesting aspect of the model, is that the mass of the lightest Higgs is around $125.5$ GeV [101, 103]. In this preliminary analysis, the hypothesis was used that only the scalars $\sigma^{0}_{2}$ and $\sigma^{\prime 0}_{2}$ get VEV. Now we will allow all the usual scalars of the MSUSY331 to have VEV, it means we want to consider that both $\sigma^{0}_{1}$ and $\sigma^{\prime 0}_{1}$ get VEV in similar conditions as presented in [54].
The goal of this work is to show that using a mechanism similar to scheme of Ma, see our Eq.(68), in MSUSY 331, we were able to remove the Majoron from the model in a similar mechanism as presented in SUSYGR. We get this nice solution, because the neutral fields belong to the sextet, ($\sigma^{\prime 0}_{1}$, $\sigma^{\prime 0}_{2}$), and to the anti-sextet, ($\sigma^{0}_{1}$), representation of the group $SU(3)_{L}$, are not necessary to generate mass for any charged fermion in the model, so their vacuum expectation values can be a few GeV, but the pseudoscalar will get large mass around ${{\cal O}}(400)$ GeV and in this way removing the non-massive Majoron from MSUSY331.
2 Minimal Supersymmetric $331$ Model.
We will introduce the following chiral superfields associated with leptons and quarks $\hat{L}_{1,2,3}$, $\hat{Q}_{1,2,3}$, $\hat{u}^{c}_{1,2,3}$, $\hat{d}^{c}_{1,2,3}$, $\hat{J}^{c}$ and $\hat{j}^{c}_{1,2}$ [101, 102]. The particle content of each chiral superfield and anti-chiral supermultiplet is presented in the Tabs.(2,3), respectively.
Chiral Superfield Fermion Scalar
$\hat{L}_{iL}=(\hat{\nu}_{i},\hat{l}_{i},\hat{l}^{c}_{i})^{T}_{L}\sim({\bf 1},{\bf 3},0)$ $L_{iL}=(\nu_{i},l_{i},l^{c}_{i})^{T}_{L}$ $\tilde{L}_{iL}=(\tilde{\nu}_{i},\tilde{l}_{i},\tilde{l}^{c}_{i})^{T}_{L}$
$\hat{Q}_{1L}=(\hat{u}_{1},\hat{d}_{1},\hat{J})^{T}_{L}\sim({\bf 3},{\bf 3},+(2/3))$ $Q_{1L}=(u_{1},d_{1},J)^{T}_{L}$ $\tilde{Q}_{1L}=(\tilde{u}_{1},\tilde{d}_{1},\tilde{J})^{T}_{L}$
Table 2: Particle content in the chiral superfields in MSUSY331 and we neglected the color indices and $i=1,2,3$. In parenthesis it appears the transformations properties under the respective factors $(SU(3)_{C},SU(3)_{L},U(1)_{N})$.
Anti-Chiral Superfield Fermion Scalar
$\hat{Q}_{\alpha L}=(\hat{d}_{\alpha},\hat{u}_{\alpha},\hat{j}_{m})^{T}_{L}\sim({\bf 3},{\bf\bar{3}},-(1/3))$ $Q_{\alpha L}=(d_{\alpha},u_{\alpha},j_{m})^{T}_{L}$ $\tilde{Q}_{\alpha L}=(\tilde{d}_{\alpha},\tilde{u}_{\alpha},\tilde{j}_{m})^{T}_{L}$
$\hat{u}^{c}_{iL}\sim({\bf\bar{3}},{\bf 1},-(2/3))$ $u^{c}_{iL}\equiv\bar{u}_{iR}$ $\tilde{u}^{c}_{iL}$
$\hat{d}^{c}_{iL}\sim({\bf\bar{3}},{\bf 1},+(1/3))$ $d^{c}_{iL}\equiv\bar{d}_{iR}$ $\tilde{d}^{c}_{iL}$
$\hat{J}^{c}_{L}\sim({\bf\bar{3}},{\bf 1},-(5/3))$ $J^{c}_{L}\equiv\bar{J}_{R}$ $\tilde{J}^{c}_{L}$
$\hat{j}^{c}_{\beta L}\sim({\bf\bar{3}},{\bf 1},+(4/3))$ $j^{c}_{\beta L}\equiv\bar{j}_{\beta R}$ $\tilde{j}^{c}_{\beta L}$
Table 3: Particle content in the anti-chiral superfields in MSUSY331 and $\alpha=2,3$, $i=1,2,3$ and $\beta=1,2$.
The usual scalars fields in the minimal, $m331$, model are defined as three triplets ($\eta,\rho,\chi$) and one anti-sextet ($S$) [88] and we can decompose those fields in
$\left(SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}\right)\supset\left(SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}\right)$ (108)
representations in the following way [51, 104, 105, 106]
  • $\left({\bf 1},{\bf 3},0\right)\rightarrow\left({\bf 1},{\bf 2},-1\right)\oplus\left({\bf 1},{\bf 1},+2\right)$
    $\eta$ $=$ $\left(\begin{array}[]{c}\Phi_{\eta}\\ \eta^{+}_{2}\end{array}\right)\sim({\bf 1},{\bf 3},0),$ (111)
    $\Phi_{\eta}$ $=$ $\left(\begin{array}[]{c}\eta^{0}\\ \eta^{-}_{1}\end{array}\right)\sim({\bf 1},{\bf 2},-1),\quad\eta^{+}_{2}\sim({\bf 1},{\bf 1},+2),$ (114)
    where the singlet $\eta^{+}_{2}$ was proposed by Zee [107]  [^14] See the Eq.(44).
  • $\left({\bf 1},{\bf 3},+1\right)\rightarrow\left({\bf 1},{\bf 2},+1\right)\oplus\left({\bf 1},{\bf 1},+4\right)$
    $\rho$ $=$ $\left(\begin{array}[]{c}\Phi_{\rho}\\ \rho^{++}\end{array}\right)\sim({\bf 1},{\bf 3},+1),$ (117)
    $\Phi_{\rho}$ $=$ $\left(\begin{array}[]{c}\rho^{+}\\ \rho^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1),\quad\rho^{++}\sim({\bf 1},{\bf 1},+4),$ (120)
  • $\left({\bf 1},{\bf 3},-1\right)\rightarrow\left({\bf 1},{\bf 2},-3\right)\oplus\left({\bf 1},{\bf 1},0\right)$
    $\chi$ $=$ $\left(\begin{array}[]{c}\Phi_{\chi}\\ \chi^{0}\end{array}\right)\sim({\bf 1},{\bf 3},-1),$ (123)
    $\Phi_{\chi}$ $=$ $\left(\begin{array}[]{c}\chi^{-}\\ \chi^{--}\end{array}\right)\sim({\bf 1},{\bf 2},-3),\quad\chi^{0}\sim({\bf 1},{\bf 1},0),$ (126)
  • $\left({\bf 1},{\bf\bar{6}},0\right)\rightarrow\left({\bf 1},{\bf 3},+2\right)\oplus\left({\bf 1},{\bf 2},+1\right)\oplus\left({\bf 1},{\bf 1},+4\right)$
    $S$ $=$ $\left(\begin{array}[]{cc}T&\frac{\Phi_{S}}{\sqrt{2}}\\ \frac{\Phi^{T}_{S}}{\sqrt{2}}&H^{--}_{2}\end{array}\right)\sim({\bf 1},{\bf{\bf\bar{6}}},0),$ (129)
    $\Phi_{S}$ $=$ $\left(\begin{array}[]{c}h^{+}_{2}\\ \sigma^{0}_{2}\end{array}\right)\sim({\bf 1},{\bf 2},+1),\quad H^{--}_{2}\sim({\bf 1},{\bf 1},+4),$ (132)
where $\Phi_{S}$ is the Higgs doublet boson of SM, see Eq.(25), and $H^{--}_{2}$ appear in the model of Babu presented in reference [108]. The triplet
$T$ $=$ $\left(\begin{array}[]{cc}\sigma^{0}_{1}&\frac{h^{+}_{1}}{\sqrt{2}}\\ \frac{h^{+}_{1}}{\sqrt{2}}&H^{--}_{1}\\ \end{array}\right)\sim({\bf 1},{\bf 3},+2),$ (135)
is the triplet $\Delta_{1}$, see Eq.(LABEL:tripletosusygelmini)  [^15] As we expect the MSUSY331 contain the SUSYGR, and the neutral scalar $\sigma^{0}_{1}$ as the scalar $\Delta^{0}_{1}$ generate Majorana mass term for the neutrinos [54].. We redefine the neutral components of the scalars defined above in the following way
$\eta^{0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\eta}+R^{0}_{\eta}+iI^{0}_{\eta}\right),\quad\rho^{0}=\frac{1}{\sqrt{2}}\left(v_{\rho}+R^{0}_{\rho}+iI^{0}_{\rho}\right),$
$\chi^{0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\chi}+R^{0}_{\chi}+iI^{0}_{\chi}\right),\quad\sigma^{0}_{1}=\frac{1}{\sqrt{2}}\left(v_{\sigma^{0}_{1}}+R^{0}_{\sigma^{0}_{1}}+iI^{0}_{\sigma^{0}_{1}}\right),$
$\sigma^{0}_{2}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\sigma^{0}_{2}}+R^{0}_{\sigma^{0}_{2}}+iI^{0}_{\sigma^{0}_{2}}\right).$ (136)
In order to implement supersymmetry, and also at same time to cancel chiral anomalies, we must to introduce the following scalars fields [101, 102]
$\eta^{\prime}$ $=$ $\left(\begin{array}[]{c}\eta^{\prime 0}\\ \eta^{\prime+}_{1}\\ \eta^{\prime-}_{2}\end{array}\right)\sim({\bf 1},{\bf\bar{3}},0),\quad\rho^{\prime}=\left(\begin{array}[]{c}\rho^{\prime-}\\ \rho^{\prime 0}\\ \rho^{\prime--}\end{array}\right)\sim({\bf 1},{\bf\bar{3}},-1),$ (143)
$\chi^{\prime}$ $=$ $\left(\begin{array}[]{c}\chi^{\prime+}\\ \chi^{\prime++}\\ \chi^{\prime 0}\end{array}\right)\sim({\bf 1},{\bf\bar{3}},+1),\quad S^{\prime}=\left(\begin{array}[]{ccc}\sigma^{\prime 0}_{1}&\frac{h^{\prime-}_{2}}{\sqrt{2}}&\frac{h^{\prime+}_{1}}{\sqrt{2}}\\ \frac{h^{\prime-}_{2}}{\sqrt{2}}&H^{\prime--}_{1}&\frac{\sigma^{\prime 0}_{2}}{\sqrt{2}}\\ \frac{h^{\prime+}_{1}}{\sqrt{2}}&\frac{\sigma^{\prime 0}_{2}}{\sqrt{2}}&H^{\prime++}_{2}\end{array}\right)\sim({\bf 1},{\bf 6},0),$ (150)
again we redefine
$\eta^{\prime 0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\eta^{\prime}}+R^{0}_{\eta^{\prime}}+iI^{0}_{\eta^{\prime}}\right),\quad\rho^{\prime 0}=\frac{1}{\sqrt{2}}\left(v_{\rho^{\prime}}+R^{0}_{\rho^{\prime}}+iI^{0}_{\rho^{\prime}}\right),$
$\chi^{\prime 0}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\chi^{\prime}}+R^{0}_{\chi^{\prime}}+iI^{0}_{\chi^{\prime}}\right),\quad\sigma^{\prime 0}_{1}=\frac{1}{\sqrt{2}}\left(v_{\sigma^{\prime 0}_{1}}+R^{0}_{\sigma^{\prime 0}_{1}}+iI^{0}_{\sigma^{\prime 0}_{1}}\right),$
$\sigma^{\prime 0}_{2}$ $=$ $\frac{1}{\sqrt{2}}\left(v_{\sigma^{\prime 0}_{2}}+R^{0}_{\sigma^{\prime 0}_{2}}+iI^{0}_{\sigma^{\prime 0}_{2}}\right).$ (152)
In the MSUSY331 we need to introduce the following three vector superfields $\hat{V}^{a}_{C}\sim({\bf 8},{\bf 1},0)$  [^16] The gluinos are the superpartner of gluons, and therefore they are in the adjoint representation of $SU(3)$, which is real., where $a=1,2,\ldots,8$, $\hat{V}^{a}\sim({\bf 1},{\bf 8},0)$, and $\hat{V}\sim({\bf 1},{\bf 1},0)$. The particle content in each vector superfield is presented in the Tab.(4). The Lagrangian of this model is presented in the references [101, 102].
${\rm{Vector\,\ Superfield}}$ ${\rm{Gauge\,\ Bosons}}$ ${\rm{Gaugino}}$ ${\rm Gauge\,\ constant}$
$\hat{V}^{a}_{C}\sim({\bf 8},{\bf 1},0)$ $g^{a}_{m}$ $\lambda_{C}^{a}$ $g_{s}$
$\hat{V}^{a}\sim({\bf 1},{\bf 8},0)$ $V^{a}_{m}$ $\lambda^{a}$ $g$
$\hat{V}\sim({\bf 1},{\bf 1},0)$ $V_{m}$ $\lambda$ $g^{\prime}$
Table 4: Particle content in the vector superfields in MSUSY331, where $a=1,2,\ldots,8$.
We denote the $SU(3)_{L}$ gauge bosons by $V^{a}_{m}$ ($a=1,2,\ldots,8$) and since
$\left({\bf 1},{\bf 8},0\right)\rightarrow\left({\bf 1},{\bf 3},0\right)\oplus\left({\bf 1},{\bf 2},3\right)\oplus\left({\bf 1},{\bf 2},-3\right)\oplus\left({\bf 1},{\bf 1},0\right),$ (153)
we get a triplet, $V^{1}_{m},V^{2}_{m},V^{3}_{m}$, and also a singlet $V^{8}_{m}$ both with $Y=0$, plus the doublet of bileptons, $\left(V^{4}_{m},V^{5}_{m}\right)$, its hypercharge is $Y=3$ and $\left(V^{6}_{m},V^{7}_{m}\right)$ and $Y=-3$. The charged gauge bosons of this model are [102]
$W^{\pm}_{m}(x)$ $=$ $-\frac{1}{\sqrt{2}}(V^{1}_{m}(x)\mp iV^{2}_{m}(x)),\,\ V^{\pm}_{m}(x)=-\frac{1}{\sqrt{2}}(V^{4}_{m}(x)\pm iV^{5}_{m}(x))$
$U^{\pm\pm}_{m}(x)$ $=$ $-\frac{1}{\sqrt{2}}(V^{6}_{m}(x)\pm iV^{7}_{m}(x)).$ (154)
The neutral gauge bosons are the photon
$A_{m}$ $=$ $\sin\theta_{W}\left(V^{3}_{m}-\sqrt{3}V^{8}_{m}\right)+\sqrt{1-4\sin^{2}\theta_{W}}V_{m},$ (155)
where $\theta_{W}$ is the Weinberg angel and we have also two massives neutral bosons defined as
$Z_{m}$ $=$ $\cos\theta_{W}V^{3}_{m}+\sqrt{3}\tan\theta_{W}\sin\theta_{W}V^{8}_{m}-\tan\theta_{W}\sqrt{1-4\sin^{2}\theta_{W}}V_{m},$
$Z^{\prime}_{m}$ $=$ $\frac{1}{\cos\theta_{W}}\left[\sqrt{1-4\sin^{2}\theta_{W}}V^{8}_{m}+\sqrt{3}\sin\theta_{W}V_{m}\right].$ (156)
In an recent article we explained the $W$-boson mass, presented by the CDF [^17] Compare with Eq.(106)., in the MSUSY331 [54]
$M^{2}_{W^{\pm}}$ $=$ $\frac{g^{2}}{4}\left[v^{2}_{SM}+v^{2}_{\sigma^{\prime}_{2}}+2\left(v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}\right)\right],$
$v^{2}_{SM}$ $=$ $v^{2}_{\eta}+v^{2}_{\rho}+v^{2}_{\sigma_{2}}+v^{2}_{\eta^{\prime}}+v^{2}_{\rho^{\prime}},$
$M^{2}_{V}$ $=$ $\frac{g^{2}}{4}\left[v^{2}_{\eta}+v^{2}_{\chi}+v^{2}_{\sigma_{2}}+v^{2}_{\eta^{\prime}}+v^{2}_{\chi^{\prime}}+v^{2}_{\sigma^{\prime}_{2}}+2\left(v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}\right)\right],$
$M^{2}_{U}$ $=$ $\frac{g^{2}}{4}\left(v^{2}_{\rho}+v^{2}_{\chi}+4v^{2}_{\sigma_{2}}+v^{2}_{\rho^{\prime}}+v^{2}_{\chi^{\prime}}+4v^{2}_{\sigma^{\prime}_{2}}\right),$
$M^{2}_{Z^{0}}$$\approx$ $\frac{1}{2}\left(\frac{g^{2}+4g^{\prime 2}}{g^{2}+3g^{\prime 2}}\right)\left[v^{2}_{SM}+v^{2}_{\sigma^{\prime}_{2}}+4\left(v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}\right)\right],$
$M^{2}_{Z^{\prime}}$$\approx$ $\frac{2}{3}(g^{2}+3g^{\prime 2})(v^{2}_{\chi}+v^{2}_{\chi^{\prime}}),$
$\rho$ $=$ $\frac{1+R}{1+R^{\prime}}\simeq 1+R-R^{\prime}=1-2\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right),$
$R$ $=$ $\frac{v^{2}_{\sigma^{\prime}_{2}}}{v^{2}_{\rm SM}}+2\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right),\,\ R^{\prime}=\frac{v^{2}_{\sigma^{\prime}_{2}}}{v^{2}_{\rm SM}}+4\left(\frac{v^{2}_{\sigma_{1}}+v^{2}_{\sigma^{\prime}_{1}}}{v^{2}_{SM}}\right).$ (157)
Those result is similar with our Eq.(106). We present some phenomenological results in A.
The superpotencial in MSUSY331 is defined as [101, 102, 103]
$W=W_{2}+W_{3}+hc,$ (158)
where each term above is defined as follows
$W_{2}$ $=$ $\sum_{i=1}^{3}\mu_{0i}(\hat{L}_{i}\hat{\eta}^{\prime})+\mu_{\eta}(\hat{\eta}\hat{\eta}^{\prime})+\mu_{\rho}(\hat{\rho}\hat{\rho}^{\prime})+\mu_{\chi}(\hat{\chi}\hat{\chi}^{\prime})+\mu_{S}Tr[(\hat{S}\hat{S}^{\prime})].$ (159)
All the coefficients in the equation above have mass dimension. The last term of the superpotential in MSUSY331 has the following form
$W_{3}$ $=$ $\sum_{i=1}^{3}\left[\sum_{j=1}^{3}\left(\sum_{k=1}^{3}\lambda_{1ijk}(\epsilon\hat{L}_{i}\hat{L}_{j}\hat{L}_{k})+\lambda_{2ij}(\epsilon\hat{L}_{i}\hat{L}_{j}\hat{\eta})+\lambda_{3ij}(\hat{L}_{i}\hat{S}\hat{L}_{j})\right)+\lambda_{4i}(\epsilon\hat{L}_{i}\hat{\chi}\hat{\rho})\right]$ (160)
$+$ $f_{1}(\epsilon\hat{\rho}\hat{\chi}\hat{\eta})+f_{2}(\hat{\chi}\hat{S}\hat{\rho})+f_{3}(\hat{\eta}\hat{S}\hat{\eta})+f_{4}\epsilon_{ijk}\epsilon_{lmn}\hat{S}_{il}\hat{S}_{jm}\hat{S}_{kn}+f^{\prime}_{1}(\epsilon\hat{\rho}^{\prime}\hat{\chi}^{\prime}\hat{\eta}^{\prime})+f^{\prime}_{2}(\hat{\chi}^{\prime}\hat{S}^{\prime}\hat{\rho}^{\prime})$
$+$ $f^{\prime}_{3}(\hat{\eta}^{\prime}\hat{S}^{\prime}\hat{\eta}^{\prime})+f^{\prime}_{4}\epsilon_{ijk}\epsilon_{lmn}\hat{S}^{\prime}_{il}\hat{S}^{\prime}_{jm}\hat{S}^{\prime}_{kn}+\sum_{i=1}^{3}\left[\kappa_{1i}(\hat{Q}_{3}\hat{\eta}^{\prime})\hat{u}^{c}_{i}+\kappa_{2i}(\hat{Q}_{3}\hat{\rho}^{\prime})\hat{d}^{c}_{i}\right]+\kappa_{3}(\hat{Q}_{3}\hat{\chi}^{\prime})\hat{J}^{c}$
$+$ $\sum_{\alpha=1}^{2}\left[\sum_{i=1}^{3}\left(\kappa_{4\alpha i}(\hat{Q}_{\alpha}\hat{\eta})\hat{d}^{c}_{i}+\kappa_{5\alpha i}(\hat{Q}_{\alpha}\hat{\rho})\hat{u}^{c}_{i}\right)+\sum_{\beta=1}^{2}\kappa_{6\alpha\beta}(\hat{Q}_{\alpha}\hat{\chi})\hat{j}^{c}_{\beta}+\sum_{i=1}^{3}\sum_{j=1}^{3}\kappa_{7\alpha ij}(\hat{Q}_{\alpha}\hat{L}_{i})\hat{d}^{c}_{j}\right]$
$+$ $\sum_{i=1}^{3}\left\{\sum_{j=1}^{3}\left[\sum_{k=1}^{3}\xi_{1ijk}\hat{d}^{c}_{i}\hat{d}^{c}_{j}\hat{u}^{c}_{k}+\sum_{\beta=1}^{2}\left(\xi_{2ij\beta}\hat{u}^{c}_{i}\hat{u}^{c}_{j}\hat{j}^{c}_{\beta}\right]+\xi_{3i\beta}\hat{d}^{c}_{i}\hat{J}^{c}\hat{j}^{c}_{\beta}\right)\right\}.$
All the coefficients in $W_{3}$ are dimensionless. Note that the scalar $S^{\prime}$ does not have any type of coupling with the fermions of the model, so the VEV of its neutral components can be a few GeV, it means our scalar field $S^{\prime}$ is similar to the scalar field $\Delta_{2}$ in the SUSYGR, see our Eq.(LABEL:suppotSUSYGR).
We want to emphasize, the gauge bosons $V^{\pm}$ and $U^{\pm\pm}$ are bileptons, see our Eq.(LABEL:lq). The term $(\hat{L}\hat{S}\hat{L})$ will introduce a term similar to Eq.(LABEL:yukawatermogelmini) then some of our scalars are also bilepton. The terms proportional for $f_{2},f_{4},f^{\prime}_{2}$ and $f^{\prime}_{4}$ [^18] They are similar to the coefficients $\lambda_{1}$ and $\lambda_{2}$ in the SUSYGR, see our Eq.(LABEL:suppotSUSYGR). violate the conservation of the quantum number
${\cal F}\equiv B+L,$ (161)
where $B$ is the baryon number while $L$ is the total lepton number  [^19] See our Eqs.(43,68) [101, 102, 103, 114]. However, if we assume the global $U(1)_{\cal F}$ symmetry, it allows us to introduce the $R$-conserving symmetry, in a similar way as we define $R$-parity in MSSM [74], defined as [109]
$R\equiv(-1)^{2S+3{\cal F}}.$ (162)
Where $S$ is the spin for a given state. The ${\cal F}$ number attribution is
$\begin{array}[]{c}{\cal F}(U^{--})={\cal F}(V^{-})=-{\cal F}(J)={\cal F}(j_{1,2})={\cal F}(\rho^{--})\\ ={\cal F}(\chi^{--})={\cal F}(\chi^{-})={\cal F}(\eta^{-}_{2})={\cal F}(\sigma_{1}^{0})=2.\end{array}$ (163)
For instance, if we allow the $\xi_{1}$ term it implies in proton decay [100]. We will show below those terms will allow us to implement the scheme of Ma, see our Eq.(68), in the MSUSY331 in the same way as we implement this mechanism in SUSYGR as we discuss after our Eq.(LABEL:suppotSUSYGR).
The soft terms to break SUSY [^20] We are not considering the Gaugino Mass terms neither sleptons and squarks, these soft terms are defined in [74]. is written as
${\cal L}^{soft}_{scalar}$ $=$ $-m^{2}_{\eta}(\eta^{\dagger}\eta)-m^{2}_{\rho}(\rho^{\dagger}\rho)-m^{2}_{\chi}(\chi^{\dagger}\chi)-m^{2}_{S}Tr[(S^{\dagger}S)]-m^{2}_{\eta^{\prime}}(\eta^{\prime\dagger}\eta^{\prime})-m^{2}_{\rho^{\prime}}(\rho^{\prime\dagger}\rho^{\prime})$ (164)
-- $m^{2}_{\chi^{\prime}}(\chi^{\prime\dagger}\chi^{\prime})-m^{2}_{S^{\prime}}Tr[(S^{\prime\dagger}S^{\prime})]+[k_{1}(\epsilon\rho\chi\eta)+k_{2}(\eta S\eta)+k_{3}(\chi S\rho)+k_{4}\epsilon\epsilon SSS$
$+$ $k^{\prime}_{1}(\epsilon\rho^{\prime}\chi^{\prime}\eta^{\prime})+k^{\prime}_{2}(\eta^{\prime}S^{\prime}\eta^{\prime})+k^{\prime}_{3}(\chi^{\prime}S^{\prime}\rho^{\prime})+k^{\prime}_{4}\epsilon\epsilon S^{\prime}S^{\prime}S^{\prime}+hc],$
$k_{1,2,3,4},k^{\prime}_{1,2,3}$ and $k^{\prime}_{4}$ has mass dimension.
The pattern of the symmetry breaking in MSUSY331 is given by the following scheme
MSUSY331 $\stackrel{{\scriptstyle{\cal L}_{soft}}}{{\longmapsto}}\mbox{SU(3)}_{C}\ \otimes\ \mbox{SU(3)}_{L}\otimes\mbox{U(1)}_{N}\stackrel{{\scriptstyle\langle\chi\rangle\langle\chi^{\prime}\rangle}}{{\longmapsto}}\mbox{SU(3)}_{C}\ \otimes\ \mbox{SU(2)}_{L}\otimes\mbox{U(1)}_{Y}$ (165)
[…]
When $\sigma_{2}$ get VEV the charged leptons obtain their masses. The fields $\sigma^{\prime}_{2},\sigma_{1}$ and $\sigma^{\prime}_{1}$ are useful to explain the $W$-boson mass presented by the CDF in the MSUSY331 in two possible ways [54]
  • 1-)
    $v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0$ GeV and $v_{\sigma_{2}}=10$ GeV and $v_{\sigma^{\prime}_{2}}=11.19$ GeV.
  • 2-)
    We can also choose
    $v_{\sigma_{1}}\neq 0,v_{\sigma^{\prime}_{1}}\neq 0,v_{\sigma^{\prime}_{2}}\neq 0.$ (166)
    In this case those VEV have to satisfy the following
    $v_{\sigma^{\prime}_{1}}=\sqrt{\frac{2\delta M^{2}_{W}}{g}-\frac{v^{2}_{\sigma^{\prime}_{2}}}{2}-v^{2}_{\sigma_{1}}},\,\ v_{\sigma^{\prime}_{2}}=8.7691841\,\ GeV,$ (167)
    some possible values is shown in our Tab.(5). We will use the following point for our numerical analyses $v_{\sigma_{1}}=4.89$ GeV, $v_{\sigma_{2}}=10$ GeV, $v_{\sigma^{\prime}_{1}}=0.5$ GeV and $v_{\sigma^{\prime}_{2}}=8.77$ GeV.
We show some details of our scalar potential in Apps.(B,C).
$v_{\sigma_{1}}$ (GeV) $v_{\sigma^{\prime}_{1}}$ (GeV)
0.5 4.89453
1.0 4.81730
1.5 4.68577
2.0 4.49515
2.5 4.2375
3.0 3.89954
3.5 3.4578
4.0 2.86468
4.5 1.98907
Table 5: Defined some values of $v_{\sigma_{1}}$ we show the numerical values for $v_{\sigma^{\prime}_{1}}$ GeV take of our study presented in [54].
3 Case 1
This case is defined by the following choice
$v_{\sigma_{1}}=v_{\sigma^{\prime}_{1}}=0\,\ GeV.$ (168)
We will use below the following set of parameters in the scalar potential
$f_{1}=1.12,\quad f_{3}=1.03,\quad f^{\prime}_{1}=f^{\prime}_{3}=10^{-6},\quad{\rm(dimensionless)}$ (169)
and
$-k_{1}=k^{\prime}_{1}=10,\;k_{3}=k^{\prime}_{3}=-100,\;-\mu_{\eta}=\mu_{\rho}=-\mu_{s}=\mu_{\chi}=1000\quad GeV,$
we also use the constraint [^21] To recover the VEV of SM given in our Eqs.(24,157).
$v^{2}_{\eta}+v^{2}_{\eta^{\prime}}+v^{2}_{\rho}+v^{2}_{\rho^{\prime}}+2(v^{2}_{\sigma^{0}_{2}}+v^{2}_{\sigma^{\prime}_{2}})=(246\;GeV)^{2}.$ (171)
Assuming that
$v_{\eta}$ $=$ $20\,\ GeV,\,\ v_{\chi}=1000\,\ GeV,\,\ v_{\sigma_{2}}=10\,\ GeV,$
$v_{\eta^{\prime}}$ $=$ $v_{\rho^{\prime}}=v_{\chi^{\prime}}=1\,\ GeV,$ (172)
the value of $v_{\rho}$ is fixed by the constraint above. We get similar values are presented in [103].
We define the following basis for the neutral scalar sector (CP even)
$\left(H_{\eta},H_{\rho},H_{\chi},H_{\sigma^{0}_{2}},H_{\eta^{\prime}},H_{\rho^{\prime}},H_{\chi^{\prime}},H_{\sigma^{\prime 0}_{2}}\right).$ (173)
All these fields are massive and we represent each mass value as $M_{H_{j}}>M_{H_{i}}$, $i=1,\cdots,8$, and with $j>i$. Our mass spectrum in (GeV) for the VEV given in our Eq.(172) is as follows
$M_{H^{0}_{1}}$$\equiv$ $M_{h^{0}}=125.577,\,\ M_{H^{0}_{2}}\equiv M_{H^{0}}=411.850,\,\ M_{H^{0}_{3}}=633.157,$
$M_{H^{0}_{4}}$ $=$ $1017.34,\,\ M_{H^{0}_{5}}=1193.52,\,\ M_{H^{0}_{6}}=1254.76,\,\ M_{H^{0}_{7}}=1375.39,$
$M_{H^{0}_{8}}$ $=$ $4791.17$ (174)
Figure 12: We show the prediction about the behaviour of the masses of lightest scalars, $M_{h^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ show in the box in several colors in case 1. $M^{exp}_{h^{0}}$ (straight line Brown) is the central experimental values given at our Eq.(28), $M^{sup}_{h^{0}}=M^{exp}_{h^{0}}+0.11$ (straight line gray) while $M^{inf}_{h^{0}}=M^{exp}_{h^{0}}-0.11$ (straight line yellow).
Figure 13: We show the prediction about the behaviour of the masses of the usual heavy scalar, $M_{H^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ show in the box in several colors in case 1. $M^{exp}_{H^{0}}$ is the lower experimental values, given in our Eq.(71), (straight line Brown).
In the neutral imaginary scalar (pseudoscalar) sector (CP odd) we define the following base
$\left(F_{\eta},F_{\rho},F_{\chi},F_{\sigma^{0}_{2}},F_{\eta^{\prime}},F_{\rho^{\prime}},F_{\chi^{\prime}},F_{\sigma^{\prime 0}_{2}}\right),$ (175)
and we have verified analytically, we have two Goldstones bosons ($G^{0}_{Z^{0}},G^{0}_{Z^{\prime 0}}$). These Goldstones bosons will give masses for the neutral gauge boson $Z^{0}$ and $Z^{\prime 0}$, respectively. The other six physical pseudoscalars have the following masses in (GeV) when we use the VEV defined in our Eq.(172), they are the following values
$M_{A^{0}_{1}}$$\equiv$ $M_{A^{0}}=411.378,\,\ M_{A^{0}_{2}}=633.183,\,\ M_{A^{0}_{3}}=1017.31,\,\ M_{A^{0}_{4}}=1193.52,$
$M_{A^{0}_{5}}$ $=$ $4791.17\displaystyle 1375.39,\,\ M_{A^{0}_{6}}=$ 4791.17 . (176)
Figure 14: We show the prediction about the behaviour of the masses of the lightest pseudoscalar, $M_{A^{0}}$, in terms of $v_{\chi}$ for several values of $v_{\chi^{\prime}}$ show in the box in several colors in case 1. $M^{exp}_{A^{0}}$ is the lower experimental values, given in our Eq.(73), (straight line Brown).
Figure 15: We show the prediction about the behaviour of the masses of the lightest pseudoscalar, $M_{A^{0}}$ and $M_{H^{0}}$, in terms of $v_{\chi}$ for $v_{\chi^{\prime}}=1$ GeV.
Possible values of the masses of scalars $h^{0},H^{0}$ and pseudoscalar $A^{0}$ in terms of $v_{\chi}$ are show in our Figs.(12,13,14), respectively. To obtain upper limits on the masses of $H^{0}$ and $A^{0}$ we draw Fig.(15), where we show $M_{H^{0}}$ and $M_{A^{0}}$ for $v_{\chi}>20$ TeV. From these figures we can say that
$900\,\ GeV$$\leq$ $v_{\chi}\leq 1000\,\ GeV,\,\ M_{h^{0}}\approx 125.20\,\ GeV;\,\ M_{H^{0}}>M_{A^{0}};$
$M_{H^{0}}$ $<$ $632\,\ GeV;\,\ M_{A^{0}}<676\,\ GeV.$ (177)
Therefore, our results for the masses of neutral scalars are within current experimental limits show in our Eqs.(28,71,73).
In our opinion, it is interesting to do a phenomenological study regarding the decays of the scalars $h^{0}$, $H^{0}$ and pseudoscalar $A^{0}$ in MSUSY331. A first series studying the decay of scalars in the 331 model without SUSY was presented in [118, 119].
It would be, also, so interesting to diagonalize these matrices in details. It means, we must to study the perturbative and unitary conditions for the potential; verify the Vacuum stability and of course the minimization condition. In our opinion, those analyses must be done first in the (m331) and then make in the MSUSY331. We have previously done a similar analysis in [^22] I am still in debt with Prof. Aahron Davidson, who introduced me to this interesting model, which should be better studied. [120].
In the Single charged Scalars we have two Goldstones bosons $G^{\pm}_{W^{\pm}}$ and $G^{\pm}_{V^{\pm}}$. They will give masses for the gauge bosons $W^{\pm}$ and $V^{\pm}$, respectively. And we also have nine massive states in two unrelated bases. The first base is defined as
$\left(\eta^{-}_{1},\rho^{-},\eta^{\prime-}_{1},\rho^{\prime-},h^{-}_{1},h^{\prime-}_{1}\right)^{T},$ (178)
we get the following masses in (GeV) for the VEV given in our Eq.(172)
$(M_{H^{\pm}_{1}})_{b1}$$\equiv$ $\left(M_{h^{\pm}}\right)_{b1}=287.399,\,\ (M_{H^{\pm}_{2}})_{b1}=970.554,\,\ (M_{H^{\pm}_{3}})_{b1}=1155.81,$
$(M_{H^{\pm}_{4}})_{b1}$ $=$ $4781.44\displaystyle 1343.41,\,\ (M_{H^{\pm}_{5}})_{b1}=$ 4781.44 . (179)
The second base is defined as
$\left(\eta^{-}_{2},\chi^{-},\eta^{\prime-}_{2},\chi^{\prime-},h^{-}_{2},h^{\prime-}_{2}\right)^{T},$ (180)
for this new basis the eigenvalues (GeV) have the following values for the VEV given in our Eq.(172)
$(M_{H^{\pm}_{1}})_{b2}$$\equiv$ $\left(M_{h^{\pm}}\right)_{b2}=357.111,\,\ (M_{H^{\pm}_{2}})_{b2}=645.823,\,\ (M_{H^{\pm}_{3}})_{b2}=931.247,$
$(M_{H^{\pm}_{4}})_{b2}$ $=$ $4781.45\displaystyle 1175.67,\,\ (M_{H^{\pm}_{5}})_{b2}=$ 4781.45 . (181)
From these data we realize that $\left(M_{h^{\pm}}\right)_{b2}$ is heavier than $\left(M_{h^{\pm}}\right)_{b1}$, which we show in more detail in our Fig.(16) for several values of $v_{\chi}$ when $v_{\chi^{\prime}}=1$ GeV. The masses values in this sector are in agreement with the actual experimental data.
Figure 16: We show the prediction about the behaviour of the masses $\left(M_{h^{\pm}}\right)_{b1}$ (straight line black) and $\left(M_{h^{\pm}}\right)_{b2}$ (straight line red) in terms of $v_{\chi}$ when $v_{\chi^{\prime}}=1$ GeV in case 1.
In the Double charged Scalars we have one Goldstone boson, $G^{\pm\pm}_{U^{\pm\pm}}$, to give mass for the gauge boson $U^{\pm\pm}$ and seven massive states, and their masses in (GeV) for the VEV given in our Eq.(172) are
$M_{H^{\pm\pm}_{1}}$$\equiv$ $M_{h^{\pm\pm}}=530.175,M_{H^{\pm\pm}_{2}}=813.436,\,\ M_{H^{\pm\pm}_{3}}=1058.42,$
$M_{H^{\pm\pm}_{4}}$ $=$ $1123.61,\,\ M_{H^{\pm\pm}_{5}}=1240.58,\,\ M_{H^{\pm\pm}_{6}}=1277.02,$
$M_{H^{\pm\pm}_{7}}$ $=$ $1396.67$ (182)
The mass of the lightest state as function of $v_{\chi}$ is shown in our Fig.(17) and they are in agreement with the actual experimental data. The mass hierarchi in this case is
$M_{h^{\pm\pm}}>\left(M_{h^{\pm}}\right)_{b2}>M_{A^{0}}>M_{H^{0}}>\left(M_{h^{\pm}}\right)_{b1}.$ (183)
Figure 17: We show the prediction about the behaviour of the masses of $M_{h^{\pm\pm}}$ in terms of $v_{\chi}$ for some values of $v_{\chi^{\prime}}$ show in the box in several colors.
4 Case 2
This case is defined by Eq.(166). Now in the sector of neutral scalar, we have two $10\times 10$. The new basis for the even CP case is
$\left(H_{\eta},H_{\rho},H_{\chi},H_{\sigma_{1}},H_{\sigma_{2}},H_{\eta^{\prime}},H_{\rho^{\prime}},H_{\chi^{\prime}},H_{\sigma^{\prime}_{1}},H_{\sigma^{\prime}_{2}}\right)^{T},$ (184)
while the base for CP odd is
$\left(F_{\eta},F_{\rho},F_{\chi},F_{\sigma_{1}},F_{\sigma_{2}},F_{\eta^{\prime}},F_{\rho^{\prime}},F_{\chi^{\prime}},F_{\sigma^{\prime}_{2}}\right)^{T}.$ (185)
For the single Charged scalars instead of having two distinct bases we have a single one, as we will present below.
If we suppose $f_{2}=f_{4}=f^{\prime}_{2}=f^{\prime}_{4}=0$, we obtain three Goldstone bosons in the CP odd sector, that is, one more Goldstone than necessary [^23] Remember $v_{\sigma^{0}_{1}}$ break $L$ symmetry as $V_{\Delta}$ do the same in Gelmini-Roncadelli scheme and $V_{\Delta_{1}},V_{\Delta_{2}}$ in SUSYGR.. This means we have, one Goldstone boson $G^{0}_{Z^{0}}$ for give mass for $Z^{0}$, another one $G^{0}_{Z^{\prime 0}}$ for give mass for $Z^{\prime 0}$ and the last one is the Majoron $M^{0}$.
We can remove this Majoron, in similar way as done in our Eq.(68), when we impose that the parameters $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ are different from zero. In other words, we see that these terms allow us to satisfactorily implement scheme of Ma, see our Eq.(68), in MSUSY331 in a similar way as implemented in SUSYGR, see our Eqs(LABEL:suppotSUSYGR,160).
We will brief present our results when $v_{\chi}=1000$ GeV and $v_{\chi^{\prime}}=1$ GeV and we also use the following numerical, in GeV, parameters
$f_{2}$ $=$ $f_{4}=1.11,\,\ f^{\prime}_{2}=f^{\prime}_{4}=10^{-6},$
$-k_{2}$ $=$ $k^{\prime}_{2}=10,\;k_{4}=k^{\prime}_{4}=-100.$ (186)
  • CP even we get and the mass spectrum in (GeV) is as follows
    $M_{H^{0}_{1}}$$\equiv$ $M_{h^{0}}=125.262,\,\ M_{H^{0}_{2}}\equiv M_{H^{0}}=422.021,\,\ M_{H^{0}_{3}}=491.822,$
    $M_{H^{0}_{4}}$ $=$ $674.681,\,\ M_{H^{0}_{5}}=1052.59,\,\ M_{H^{0}_{6}}=1200.45,\,\ M_{H^{0}_{7}}=1281.98,$
    $M_{H^{0}_{8}}$ $=$ $1369.16,\,\ M_{H^{0}_{9}}=1411.46,\,\ M_{H^{0}_{10}}=4777.62;$ (187)
  • CP odd the base have the following masses, with the same parameters as before, in (GeV)
    $M_{A^{0}_{1}}$$\equiv$ $M_{A^{0}}=420.963,\,\ M_{A^{0}_{2}}=495.397,\,\ M_{A^{0}_{3}}=674.74,$
    $M_{A^{0}_{4}}$ $=$ $1052.56,\,\ M_{A^{0}_{5}}=1200.45,\,\ M_{A^{0}_{6}}=1369.17,\,\ M_{A^{0}_{7}}=1411.42,$
    $M_{A^{0}_{8}}$ $=$ $4777.61;$ (188)
  • Single charged Scalars we get the following masses, in (GeV)
    $M_{H^{\pm}_{1}}$$\equiv$ $M_{h^{\pm}}=299.376,\,\ M_{H^{\pm}_{2}}=334.292,\,\ M_{H^{\pm}_{3}}=440.198,$
    $M_{H^{\pm}_{4}}$ $=$ $632.72,\,\ M_{H^{\pm}_{5}}=931.934,\,\ M_{H^{\pm}_{6}}=967.829,\,\ M_{H^{\pm}_{7}}=1307.3,$
    $M_{H^{\pm}_{8}}$ $=$ $1371.22,\,\ M_{H^{\pm}_{9}}=4731.05,\,\ M_{H^{\pm}_{10}}=4760.51;$ (189)
  • Double charged Scalars we get the following masses, in (GeV)
    $M_{H^{\pm\pm}_{1}}$$\equiv$ $M_{h^{\pm\pm}}=526.523,\,\ M_{H^{\pm\pm}_{2}}=537.62,\,\ M_{H^{\pm\pm}_{3}}=804.524,$
    $M_{H^{\pm\pm}_{4}}$ $=$ $989.578,\,\ M_{H^{\pm\pm}_{5}}=999.689,\,\ M_{H^{\pm\pm}_{6}}=1247.99,$
    $M_{H^{\pm\pm}_{7}}$ $=$ $1311.01$ (190)
We also get the results show in our Tabs.(6,7), where we notice the masses for $M_{h^{0}}$, $M_{H^{0}}$, $M_{A^{0}}$, $M_{h^{\pm}}$ and $M_{h^{\pm\pm}}$ are in agreement with the actual experimental values and we can conclude
  • $M_{h^{0}}\sim 125.5$ GeV;
  • $M_{H^{0}}>400$ GeV; $M_{A^{0}}>400$ GeV;
  • $M_{h^{\pm}}>M_{h^{\pm\pm}}>M_{H^{0}}>M_{A^{0}}$.
$v_{\chi^{\prime}}$ (GeV) $M_{h^{0}}$ (GeV) $M_{H^{0}}$ (GeV) $M_{A^{0}}$ (GeV) $M_{h^{\pm}}$ (GeV)
15.30 125.596 422.381 422.273 813.965
15.35 125.483 422.373 422.271 813.930
15.40 125.396 422.366 422.269 813.896
15.45 125.309 422.358 422.267 813.861
15.50 125.220 422.350 422.264 813.826
Table 6: The masses of $M_{h^{0}},M_{H^{0}},M_{A^{0}}$ and $M_{h^{\pm}}$ for $v_{\chi}=10$ TeV and some values of $v_{\chi^{\prime}}$ and under this suppositions we get $M_{h^{\pm\pm}}=527.140$ GeV.
$v_{\chi}$ (GeV) $M_{h^{0}}$ (GeV) $M_{H^{0}}$ (GeV) $M_{A^{0}}$ (GeV)
8500 124.871 422.349 422.223
9000 124.996 422.349 422.245
9500 125.112 422.350 422.255
10000 125.220 422.350 422.264
10500 125.323 422.351 422.271
Table 7: The masses of $M_{h^{0}},M_{H^{0}},M_{A^{0}}$ and for $v_{\chi^{\prime}}=15.5$ GeV and some values of $v_{\chi^{\prime}}$ and under this suppositions we get $M_{h^{\pm}}813.826$ GeV and $M_{h^{\pm\pm}}=527.140$ GeV.
5 Conclusions
We have studied the scalar potential of supersymmetric $331$ model with all the VEV from the sextet and anti-sextet are not null and we show our lightest scalar fields has mass near the value $M_{h^{0}}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(28). We also show that in MSUSY331, we have $M_{A^{0}},M_{H^{0}}>400$ GeV, $M_{h^{\pm}}>181$ GeV and $M_{h^{\pm\pm}}>445$ GeV. Regarding which is heavier, $H^{0}$ or $A^{0}$, the answer depends on the condition. We show that in case 1 we have $M_{A^{0}}>M_{H^{0}}$, see our Fig.(15), and in case 2 the opposite, that is, $M_{H^{0}}>M_{A^{0}}$ as we show in our Tab.(7).
Acknowledgments
We would like thanks V. Pleitez for useful discussions above $331$ models. We also to thanks IFT for the nice hospitality during my several visit to perform my studies about the $331$ models and also for done this article.
Appendix A Some phenomenological Results in MSUSY331.
The interaction between the charged bosons with the leptons are given by [102, 109]
${\cal L}_{l}^{CC}$ $=$ $-\frac{g}{\sqrt{2}}\sum_{l}\left(\bar{\nu}_{lL}\gamma^{m}V_{\rm PMNS}l_{L}W^{+}_{m}+\bar{l}^{c}_{L}\gamma^{m}U_{V}\nu_{lL}V^{+}_{m}+\bar{l}^{c}_{L}\gamma^{m}U_{U}l_{L}U^{++}_{m}+hc\right).$
The $V_{\rm PMNS}$ is the Pontecorvo-Maki-Nakagawa-Sakata mixing matrix. There are new mixing matrices given by $U_{V}$ and $U_{U}$. The bosons $U^{--}$ and $V^{-}$ are called bileptons because they couple to two leptons; thus they have two units of lepton number, it means $L=L_{e}+L_{\mu}+L_{\tau}=2$. This model does not conserve separate family lepton number, $L_{e}$, $L_{\mu}$ and $L_{\tau}$ but only the total lepton number $L$ is conserved. We have already showed that in the Møller scattering and in muon-muon scattering we can show that left-right asymmetries $A_{RL}(ll)$ are very sensitive to a doubly charged vector bilepton resonance but they are insensitive to scalar ones [110, 111, 112].
Similarly, we have the neutral currents coupled to both $Z^{0}$ and $Z^{\prime 0}$ massive vector bosons, according to the Lagrangian
${\cal L}_{\nu}^{NC}=-\frac{g}{2}\frac{M_{Z^{0}}}{M_{W^{\pm}}}\bar{\nu}_{lL}\gamma^{m}\nu_{lL}\left[Z^{0}_{m}-\frac{1}{\sqrt{3}}\frac{1}{\sqrt{h(t)}}Z^{\prime 0}_{m}\right],$ (192)
with
$h(t)=1+4t^{2},$ (193)
for neutrinos and
${\cal L}_{l}^{NC}=-\frac{g}{4}\frac{M_{Z^{0}}}{M_{W^{\pm}}}\left[\bar{l}\gamma^{m}(v_{l}+a_{l}\gamma^{5})lZ_{m}+\bar{l}\gamma^{m}(v^{\prime}_{l}+a^{\prime}_{l}\gamma^{5})lZ^{\prime 0}_{m}\right],$ (194)
for the charged leptons, where we have defined
$\begin{array}[]{clccc}v_{l}=&-\frac{1}{h(t)},&a_{l}=&1,&\\ v^{\prime}_{l}=&-\sqrt{\frac{3}{h(t)}},&a^{\prime}_{l}=&\frac{v^{\prime}_{l}}{3}.&\end{array}$
We can use muon collider to discover the new neutral $Z^{\prime}$ boson using the reaction $\mu e\to\mu e$ it was show at [111, 113] that $A_{RL}(\mu e)$ asymmetry is considerably enhanced, some phenomenological analyses was presented by references [101, 102, 109].
For some phenomenological analyses, it is useful to define the following parameters [115, 116]  [^24] We want to perform similar constraints as presented in [115, 116] in the MSUSY331.
$\xi_{WW^{\prime}}=\left(\frac{M_{W}}{M_{W^{\prime}}}\right)^{2},\,\ \xi_{ZZ^{\prime}}=\left(\frac{M_{Z}}{M_{Z^{\prime}}}\right)^{2}.$ (196)
We show some values of these parameters in MSUSY331 in our Tabs.(8,9).
$v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
Table 8: The parameters $\xi_{WW^{\prime}}$ and $\xi_{ZZ^{\prime}}$, see Eq.(196), as function of $v_{\chi}$ in GeV and we take $v_{\chi^{\prime}}=1000$ GeV and the others VEV are given in Eqs.(63,64) and Eq.(73) [54].
$v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
Table 9: The parameters $\xi_{WW^{\prime}}$ and $\xi_{ZZ^{\prime}}$, see Eq.(196), as function of $v_{\chi}$ in GeV and we take $v_{\chi^{\prime}}=1000$ GeV and the others VEV are given in Eqs.(63,64) and Eq.(73) [54].
We would like to highlight, the new exotic quarks, $J$, $j_{1}$ and $j_{2}$, may be discovered by the Large Hadron Collider (LHC) thought $pp$ collisions, via the following subprocess [^25] I want to thank Alexander S. Belyaev, who brought this process to my attention at the end of my PhD studies at IFT-Unesp, but unfortunately we were unable to publish this study together.
$g+d\rightarrow U^{--}+J,\,\ g+u\rightarrow U^{--}+j_{\alpha},$ (197)
its signature is $llXX$ and it can be detected at LHC if they really exist in nature [109, 117]. Another interesting process, with the same previous signature, to prove this model is
$e^{-}+e^{-}\rightarrow\tilde{\chi}^{--}+\tilde{\chi}^{0},$ (198)
it was presented by [^26] I want to thank M. Capdequi-Peyranère, because he taught me a lot about the phenomenology of supersymmetric models in 2000 when I started studying this model. [101, 102, 109].
Appendix B Review Scalar Potential at m331
The most general scalar potential involving triplets and the anti-sextet is [106]
$V(\eta,\rho,\chi,S)$ $=$ $V^{(2)}+V^{(3)}+V^{(4a)}+V^{(4b)}+V^{(4c)}+V^{(4d)}+V^{(4e)},$ (199)
where
$V^{(2)}$ $=$ $\mu_{1}^{2}(\eta^{\dagger}\eta)+\mu_{2}^{2}(\rho^{\dagger}\rho)+\mu_{3}^{2}(\chi^{\dagger}\chi)+\mu_{4}^{2}Tr[(S^{\dagger}S)],$
$V^{(3)}$ $=$ $\frac{f_{1}}{3!}\varepsilon_{ijk}\eta_{i}\rho_{j}\chi_{k}+f_{2}\rho_{i}\chi_{j}S^{\dagger ij}+f_{3}\eta_{i}\eta_{j}S^{\dagger ij}+\frac{f_{4}}{3!}\epsilon_{ijk}\epsilon_{lmn}S_{il}S_{jm}S_{kn}+hc,$
$V^{(4a)}$ $=$ $a_{1}(\eta^{\dagger}\eta)^{2}+a_{2}(\rho^{\dagger}\rho)^{2}+a_{3}(\chi^{\dagger}\chi)^{2}+(\chi^{\dagger}\chi)\left[a_{4}(\eta^{\dagger}\eta)+a_{5}(\rho^{\dagger}\rho)\right]+a_{6}(\eta^{\dagger}\eta)(\rho^{\dagger}\rho)$
$+$ $a_{7}(\chi^{\dagger}\eta)(\eta^{\dagger}\chi)+a_{8}(\chi^{\dagger}\rho)(\rho^{\dagger}\chi)+a_{9}(\eta^{\dagger}\rho)(\rho^{\dagger}\eta)+\left[a_{10}(\chi^{\dagger}\eta)(\rho^{\dagger}\eta)+hc\right],$
$V^{(4b)}$ $=$ $b_{1}(\chi^{\dagger}S)(\hat{\chi}\eta)+b_{2}(\rho^{\dagger}S)(\hat{\rho}\eta)+b_{3}(\eta^{\dagger}S)\left[(\hat{\chi}\rho)-(\hat{\rho}\chi)\right]+hc,$
$V^{(4c)}$ $=$ $c_{1}Tr[(\hat{\eta}S)(\hat{\eta}S)]+c_{2}Tr[(\hat{\rho}S)(\hat{\rho}S)]+hc,$
$V^{(4d)}$ $=$ $d_{1}(\chi^{\dagger}\chi)Tr[(S^{\dagger}S)]+d_{2}Tr[(\chi^{\dagger}S)(S^{\dagger}\chi)]+d_{3}(\eta^{\dagger}\eta)Tr[(S^{\dagger}S)]+d_{4}Tr[(\eta^{\dagger}S)(S^{\dagger}\eta)]$
$+$ $d_{5}(\rho^{\dagger}\rho)Tr[(S^{\dagger}S)]+d_{6}Tr[(\rho^{\dagger}S)(S^{\dagger}\rho)],$
$V^{(4e)}$ $=$ $e_{1}(Tr[(S^{\dagger}S)])^{2}+e_{2}Tr[(S^{\dagger}S)(S^{\dagger}S)],$ (200)
and we have defined the $V^{(4b)}$ and $V^{(4c)}$ terms
$\hat{x}_{ij}\equiv\epsilon_{ijk}x_{k},$ (201)
with $x=\eta,\rho,\chi$. The constants $f_{i}$, $i=1,2,3$ and $4$ have dimension of mass.
Appendix C Construction Scalar Potential
To get the scalar potential of our model we have to eliminate the auxiliarly fields $F$ and $D$ that appear in our model. We are going to pick up the $F$ and $D$- terms we get
${\cal L}^{Gauge}_{D}$ $=$ $\frac{1}{2}D^{a}D^{a}+\frac{1}{2}DD\,\ ,$
${\cal L}^{Scalar}_{F}$ $=$ $|F_{\eta}|^{2}+|F_{\rho}|^{2}+|F_{\chi}|^{2}+|F_{S}|^{2}+|F_{\eta^{\prime}}|^{2}+|F_{\rho^{\prime}}|^{2}+|F_{\chi^{\prime}}|^{2}+|F_{S^{\prime}}|^{2},$
${\cal L}^{Scalar}_{D}$ $=$ $\frac{g}{2}\left[\bar{\eta}\lambda^{a}\eta+\bar{\rho}\lambda^{a}\rho+\bar{\chi}\lambda^{a}\chi+\bar{S}\lambda^{a}S-\bar{\eta}^{\prime}\lambda^{*a}\eta^{\prime}-\bar{\rho}^{\prime}\lambda^{*a}\rho^{\prime}-\bar{\chi}^{\prime}\lambda^{*a}\chi^{\prime}-\bar{S}^{\prime}\lambda^{*a}S^{\prime}\right]D^{a}$
$+$ $\frac{g^{\prime}}{2}\left[\bar{\rho}\rho-\bar{\chi}\chi-\bar{\rho}^{\prime}\rho^{\prime}+\bar{\chi}^{\prime}\chi^{\prime}\right]D,$
${\cal L}^{W2}_{F}$ $=$ $\mu_{\eta}(\eta F_{\eta^{\prime}}+\eta^{\prime}F_{\eta})+\mu_{\rho}(\rho F_{\rho^{\prime}}+\rho^{\prime}F_{\rho})+\mu_{\chi}(\chi F_{\chi^{\prime}}+\chi^{\prime}F_{\chi})+hc,$
${\cal L}^{W3}_{F}$ $=$ $f_{1}\epsilon(F_{\rho}\chi\eta+\rho F_{\chi}\eta+\rho\chi F_{\eta})+f_{2}(2F_{\eta}\eta S+\eta\eta F_{S})+f_{3}(F_{\rho}S\chi+\rho F_{S}\chi+\rho SF_{\chi})$
$+$ $3f_{4}\epsilon\epsilon SSF_{S}+f^{\prime}_{1}\epsilon(F_{\rho^{\prime}}\chi^{\prime}\eta^{\prime}+\rho^{\prime}F_{\chi^{\prime}}\eta^{\prime}+\rho^{\prime}\chi^{\prime}F_{\eta^{\prime}})+f^{\prime}_{2}(2F_{\eta^{\prime}}\eta^{\prime}S^{\prime}+\eta^{\prime}\eta^{\prime}F_{S^{\prime}})$
$+$ $f^{\prime}_{3}(F_{\rho^{\prime}}S^{\prime}\chi^{\prime}+\rho^{\prime}F_{S^{\prime}}\chi^{\prime}+\rho^{\prime}S^{\prime}F_{\chi^{\prime}})+3f^{\prime}_{4}\epsilon\epsilon S^{\prime}S^{\prime}F_{S^{\prime}}+hc.$
From the equation described above we can construct
${\cal L}_{F}$ $=$ ${\cal L}^{Scalar}_{F}+{\cal L}^{W2}_{F}+{\cal L}^{W3}_{F}$
$=$ $|F_{\eta}|^{2}+|F_{\rho}|^{2}+|F_{\chi}|^{2}+|F_{\eta^{\prime}}|^{2}+|F_{\rho^{\prime}}|^{2}+|F_{\chi^{\prime}}|^{2}$
$+$ $\mu_{\eta}(\eta F_{\eta^{\prime}}+\eta^{\prime}F_{\eta})+\mu_{\rho}(\rho F_{\rho^{\prime}}+\rho^{\prime}F_{\rho})+\mu_{\chi}(\chi F_{\chi^{\prime}}+\chi^{\prime}F_{\chi})$
$+$ $f_{1}\epsilon(F_{\rho}\chi\eta+\rho F_{\chi}\eta+\rho\chi F_{\eta})+f_{2}(2F_{\eta}\eta S+\eta\eta F_{S})+f_{3}(F_{\rho}S\chi+\rho F_{S}\chi+\rho SF_{\chi})$
$+$ $3f_{4}\epsilon\epsilon SSF_{S}+f^{\prime}_{1}\epsilon(F_{\rho^{\prime}}\chi^{\prime}\eta^{\prime}+\rho^{\prime}F_{\chi^{\prime}}\eta^{\prime}+\rho^{\prime}\chi^{\prime}F_{\eta^{\prime}})+f^{\prime}_{2}(2F_{\eta^{\prime}}\eta^{\prime}S^{\prime}+\eta^{\prime}\eta^{\prime}F_{S^{\prime}})$
$+$ $f^{\prime}_{3}(F_{\rho^{\prime}}S^{\prime}\chi^{\prime}+\rho^{\prime}F_{S^{\prime}}\chi^{\prime}+\rho^{\prime}S^{\prime}F_{\chi^{\prime}})+3f^{\prime}_{4}\epsilon\epsilon S^{\prime}S^{\prime}F_{S^{\prime}}+hc\,\ ,$
${\cal L}_{D}$ $=$ ${\cal L}^{Gauge}_{D}+{\cal L}^{Scalar}_{D}$ (203)
$=$ $\frac{1}{2}D^{a}D^{a}+\frac{1}{2}DD+\frac{g}{2}\left[\bar{\eta}\lambda^{a}\eta+\bar{\rho}\lambda^{a}\rho+\bar{\chi}\lambda^{a}\chi-\bar{\eta}^{\prime}\lambda^{*a}\eta^{\prime}-\bar{\rho}^{\prime}\lambda^{*a}\rho^{\prime}\right.$
-- $\left.\bar{\chi}^{\prime}\lambda^{*a}\chi^{\prime}\right]D^{a}+\frac{g^{\prime}}{2}\left[\bar{\rho}\rho-\bar{\chi}\chi-\bar{\rho}^{\prime}\rho^{\prime}+\bar{\chi}^{\prime}\chi^{\prime}\right]D.$
We will now show that these fields can be eliminated through the Euler-Lagrange equations
$\frac{\partial{\cal L}}{\partial\phi}-\partial_{m}\frac{\partial{\cal L}}{\partial(\partial_{m}\phi)}=0\,\ ,$ (204)
where $\phi=\eta,\rho,\chi,S,\eta^{\prime},\rho^{\prime},\chi^{\prime},S^{\prime}$. Formally auxiliary fields are defined as fields having no kintetic terms. Thus, this definition immediately yields that the Euler-Lagrange equations for auxiliary fields simplify to $\frac{\partial{\cal L}}{\partial\phi}=0$.
$V_{D}$ $=$ $-{\cal L}_{D}=\frac{1}{2}\left(D^{a}D^{a}+DD\right)$ (205)
$=$ $\frac{g^{\prime 2}}{2}(\rho^{\dagger}\rho-\rho^{\prime\dagger}\rho^{\prime}-\chi^{\dagger}\chi+\chi^{\prime\dagger}\chi^{\prime})^{2}+\frac{g^{2}}{8}\sum_{i,j}\left(\eta^{\dagger}_{i}\lambda^{a}_{ij}\eta_{j}+\rho^{\dagger}_{i}\lambda^{a}_{ij}\rho_{j}+\chi^{\dagger}_{i}\lambda^{a}_{ij}\chi_{j}+S^{\dagger}_{ij}\lambda^{a}_{jk}S_{kl}\right.$
-- $\left.\eta^{\prime\dagger}_{i}\lambda^{*a}_{ij}\eta^{\prime}_{j}-\rho^{\prime\dagger}_{i}\lambda^{*a}_{ij}\rho^{\prime}_{j}-\chi^{\prime\dagger}_{i}\lambda^{*a}_{ij}\chi^{\prime}_{j}-S^{\prime\dagger}_{ij}\lambda^{*a}_{jk}S^{\prime}_{kl}\right)^{2},$
We can use the following relations
$\lambda^{a}_{ij}\lambda^{a}_{kl}=\frac{(-2)}{3}\delta_{ij}\delta_{kl}+2\delta{il}\delta_{jk}.$ (206)
Then
$\eta^{\dagger}_{i}\eta_{j}\eta^{\dagger}_{k}\eta_{l}\lambda^{a}_{ij}\lambda^{a}_{kl}=-\frac{2}{3}\left(\eta^{\dagger}\eta\right)\left(\eta^{\dagger}\eta\right)+2\left(\eta^{\dagger}\eta\right)\left(\eta^{\dagger}\eta\right)=\frac{4}{3}\left(\eta^{\dagger}\eta\right)^{2},$ (207)
in similar way
$\rho^{\dagger}_{i}\rho_{j}\rho^{\dagger}_{k}\rho_{l}\lambda^{a}_{ij}\lambda^{a}_{kl}$ $=$ $\frac{4}{3}\left(\rho^{\dagger}\rho\right)^{2},$
$\chi^{\dagger}_{i}\chi_{j}\chi^{\dagger}_{k}\chi_{l}\lambda^{a}_{ij}\lambda^{a}_{kl}$ $=$ $\frac{4}{3}\left(\chi^{\dagger}\chi\right)^{2},$
$S^{\dagger}_{il}S_{mj}S^{\dagger}_{kn}S_{op}\lambda^{a}_{lm}\lambda^{a}_{no}$ $=$ $-\frac{2}{3}{\mbox{T}r}\left[\left(S^{\dagger}S\right)\left(S^{\dagger}S\right)\right]+2{\mbox{T}r}\left[(S^{\dagger}S)(S^{\dagger}S)\right].$ (208)
By another hand
$\eta^{\dagger}_{i}\eta_{j}\rho^{\dagger}_{k}\rho_{l}\lambda^{a}_{ij}\lambda^{a}_{kl}$ $=$ $-\frac{2}{3}\left(\eta^{\dagger}\eta\right)\left(\rho^{\dagger}\rho\right)+2\left(\rho^{\dagger}\eta\right)\left(\eta^{\dagger}\rho\right),$
$\eta^{\dagger}_{i}\eta_{j}\chi^{\dagger}_{k}\chi_{l}\lambda^{a}_{ij}\lambda^{a}_{kl}$ $=$ $-\frac{2}{3}\left(\eta^{\dagger}\eta\right)\left(\chi^{\dagger}\chi\right)+2\left(\chi^{\dagger}\eta\right)\left(\eta^{\dagger}\chi\right),$
$\rho^{\dagger}_{i}\rho_{j}\chi^{\dagger}_{k}\chi_{l}\lambda^{a}_{ij}\lambda^{a}_{kl}$ $=$ $-\frac{2}{3}\left(\rho^{\dagger}\rho\right)\left(\chi^{\dagger}\chi\right)+2\left(\chi^{\dagger}\rho\right)\left(\rho^{\dagger}\chi\right),$ (209)
We can rewrite $V_{D}$ in the following way
$V_{D}$ $=$ $\frac{g^{\prime 2}}{2}\left[(\rho^{\dagger}\rho)^{2}+(\chi^{\dagger}\chi)^{2}-2(\rho^{\dagger}\rho)(\chi^{\dagger}\chi)+\ldots\right]$ (210)
$+$ $\frac{g^{2}}{8}\left\{\frac{(-4)}{3}\left[(\eta^{\dagger}\eta)^{2}+(\rho^{\dagger}\rho)^{2}+(\chi^{\dagger}\chi)^{2}+(\eta^{\dagger}\eta)(\rho^{\dagger}\rho)+(\eta^{\dagger}\eta)(\chi^{\dagger}\chi)\right.\right.$
$+$ $\left.\left.(\eta^{\dagger}\eta)Tr[(S^{\dagger}S)]+(\rho^{\dagger}\rho)(\chi^{\dagger}\chi)+(\rho^{\dagger}\rho)Tr[(S^{\dagger}S)]+(\chi^{\dagger}\chi)Tr[(S^{\dagger}S)]\right]\right.$
$+$ $\left.4\left[(\eta^{\dagger}\rho)(\rho^{\dagger}\eta)+(\eta^{\dagger}\chi)(\chi^{\dagger}\eta)+Tr[(S^{\dagger}\eta)(\eta^{\dagger}S)]+(\rho^{\dagger}\chi)(\chi^{\dagger}\rho)\right.\right.$
$+$ $\left.\left.Tr[(S^{\dagger}\rho)(\rho^{\dagger}S)]+Tr[(S^{\dagger}\chi)(\chi^{\dagger}S)]\right]-\frac{2}{3}\left(Tr[(S^{\dagger}S)]\right)^{2}\right.$
$+$ $\left.2Tr[(S^{\dagger}S)(S^{\dagger}S)]+\ldots\right\}.$
where $\ldots$ include the new fields due the SUSY algebra. As we want to compare our potential with the m331 we omitt those terms.
$V_{F}$ $=$ $-{\cal L}_{F}=\sum_{m}F^{*}_{m}F_{m}$
$=$ $\sum_{i,j,k}\left[\left|\mu_{\eta}\eta^{\prime}_{i}+f_{1}\epsilon_{ijk}\rho_{j}\chi_{k}+2f_{2}\eta_{i}S_{ij}\right|^{2}+\left|\mu_{\rho}\rho^{\prime}_{i}+f_{1}\epsilon_{ijk}\chi_{j}\eta_{k}+f_{3}\chi_{i}S_{ij}\right|^{2}\right.$
$+$ $\left.\left|\mu_{\chi}\chi^{\prime}_{i}+f_{1}\epsilon_{ijk}\rho_{j}\eta_{k}+f_{3}\rho_{i}S_{ij}\right|^{2}+\left|\mu_{S}S^{\prime}_{ij}+f_{2}\eta_{i}\eta_{j}+f_{3}\chi_{i}\rho_{j}+3f_{4}\epsilon_{ilk}\epsilon_{jmn}S_{lm}S_{kn}\right|^{2}\right.$
$+$ $\left.\left|\mu_{\eta}\eta_{i}+f^{\prime}_{1}\epsilon_{ijk}\rho^{\prime}_{j}\chi^{\prime}_{k}+2f^{\prime}_{2}\eta^{\prime}_{i}S^{\prime}_{ij}\right|^{2}+\left|\mu_{\rho}\rho_{i}+f^{\prime}_{1}\epsilon_{ijk}\chi^{\prime}_{j}\eta^{\prime}_{k}+f^{\prime}_{3}\chi^{\prime}_{i}S^{\prime}_{ij}\right|^{2}\right.$
$+$ $\left.\left|\mu_{\chi}\chi_{i}+f^{\prime}_{1}\epsilon_{ijk}\rho^{\prime}_{j}\eta^{\prime}_{k}+f^{\prime}_{3}\rho^{\prime}_{i}S^{\prime}_{ij}\right|^{2}+\left|\mu_{S}S_{ij}+f^{\prime}_{2}\eta^{\prime}_{i}\eta^{\prime}_{j}+f^{\prime}_{3}\chi^{\prime}_{i}\rho^{\prime}_{j}+3f^{\prime}_{4}\epsilon_{ilk}\epsilon_{jmn}S^{\prime}_{lm}S^{\prime}_{kn}\right|^{2}\right]$
We can use the following relation
$\epsilon_{ijk}\epsilon_{kmn}=\delta_{im}\delta_{jn}-\delta_{in}\delta_{jm}.$ (212)
As we want to get the M331 we will make [^27] To satisfy ${\cal F}$ defined in Eq.(161).
$f_{2}=f_{4}=0=f^{\prime}_{2}=f^{\prime}_{4},$ (213)
then we get
$\left|\mu_{\eta}\eta_{i}+\ldots\right|^{2}$ $=$ $\mu_{\eta}(\eta^{\dagger}\eta),\,\ \left|f_{1}\epsilon_{ijk}\rho_{j}\chi_{k}+\ldots\right|^{2}=f^{2}_{1}\left[(\rho^{\dagger}\rho)(\chi^{\dagger}\chi)-(\rho^{\dagger}\chi)(\chi^{\dagger}\rho)\right],$
$\left|\mu_{\rho}\rho_{i}+\ldots\right|^{2}$ $=$ $\mu^{2}_{\rho}(\rho^{\dagger}\rho),\,\ \left|\mu_{\chi}\chi_{i}+\ldots\right|^{2}=\mu^{2}_{\chi}(\chi^{\dagger}\chi),$
$\left|f_{1}\epsilon_{ijk}\chi_{j}\eta_{k}+f_{3}\chi_{i}S_{ij}+\ldots\right|^{2}$ $=$ $f^{2}_{1}\left[(\chi^{\dagger}\chi)(\eta^{\dagger}\eta)-(\chi^{\dagger}\eta)(\eta^{\dagger}\chi)\right]+f_{1}f_{3}\left[\epsilon(\chi^{\dagger}\eta^{\dagger})(\chi S)+hc\right]$
$+$ $f^{2}_{3}{\mbox{T}r}\left[(S^{\dagger}\chi)(\chi^{\dagger}S)\right],$
$\left|f_{1}\epsilon_{ijk}\rho_{j}\eta_{k}+f_{3}\rho_{i}S_{ij}+\ldots\right|^{2}$ $=$ $f^{2}_{1}\left[(\rho^{\dagger}\rho)(\eta^{\dagger}\eta)-(\rho^{\dagger}\eta)(\eta^{\dagger}\rho)\right]+f_{1}f_{3}\left[\epsilon(\rho^{\dagger}\eta^{\dagger})(\rho S)+hc\right]$
$+$ $f^{2}_{3}{\mbox{T}r}\left[(S^{\dagger}\rho)(\rho^{\dagger}S)\right],$
$\left|\mu_{S}S_{ij}+\ldots\right|^{2}$ $=$ $\mu^{2}_{S}(Tr[(S^{\dagger}S)])^{2},$
$\left|f_{1}\chi_{i}\rho_{j}+\ldots\right|^{2}$ $=$ $f^{2}_{1}\left[(\rho^{\dagger}\rho)(\chi^{\dagger}\chi)-(\rho^{\dagger}\chi)(\chi^{\dagger}\rho)\right].$
Therefore
$V_{F}$ $=$ $\mu^{2}_{\eta}(\eta^{\dagger}\eta)^{2}+\mu^{2}_{\rho}(\rho^{\dagger}\rho)^{2}+\mu^{2}_{\chi}(\chi^{\dagger}\chi)^{2}+\mu^{2}_{S}(Tr[(S^{\dagger}S)])^{2}$ (215)
$+$ $f^{2}_{1}\left[(\eta^{\dagger}\eta)(\rho^{\dagger}\rho)-(\rho^{\dagger}\eta)(\eta^{\dagger}\rho)+(\eta^{\dagger}\eta)(\chi^{\dagger}\chi)\right.$
-- $\left.(\chi^{\dagger}\eta)(\eta^{\dagger}\chi)+(\rho^{\dagger}\rho)(\chi^{\dagger}\chi)-(\rho^{\dagger}\chi)(\chi^{\dagger}\rho)\right]+f^{2}_{3}\left\{Tr[(S^{\dagger}\rho)(\rho^{\dagger}S)]+Tr[(S^{\dagger}\chi)(\chi^{\dagger}S)]\right\}$
$+$ $\left\{f_{1}f_{3}\left[\epsilon(\rho^{\dagger}\eta^{\dagger})(\rho S)+\epsilon(\chi^{\dagger}\eta^{\dagger})(\chi S)\right]+hc\right\}+\ldots,$
The scalar potential in this model is
$V_{scalar}$ $=$ $m^{2}_{\eta}(\eta^{\dagger}\eta)+m^{2}_{\rho}(\rho^{\dagger}\rho)+m^{2}_{\chi}(\chi^{\dagger}\chi)+m^{2}_{S}Tr[(S^{\dagger}S)]-[k_{1}\epsilon\rho\chi\eta+k_{3}\chi\rho S^{\dagger}+hc]$ (216)
$+$ $\left(\mu^{2}_{\eta}-\frac{g^{2}}{6}\right)(\eta^{\dagger}\eta)^{2}+\left(\mu^{2}_{\rho}+\frac{g^{\prime 2}}{2}-\frac{g^{2}}{6}\right)(\rho^{\dagger}\rho)^{2}+\left(\mu^{2}_{\chi}+\frac{g^{\prime 2}}{2}-\frac{g^{2}}{6}\right)(\chi^{\dagger}\chi)^{2}$
$+$ $\left(f^{2}_{1}-\frac{g^{2}}{6}\right)(\eta^{\dagger}\eta)(\rho^{\dagger}\rho)+\left(f^{2}_{1}-\frac{g^{2}}{6}\right)(\eta^{\dagger}\eta)(\chi^{\dagger}\chi)+\left(f^{2}_{1}-g^{\prime 2}-\frac{g^{2}}{6}\right)(\rho^{\dagger}\rho)(\chi^{\dagger}\chi)$
$+$ $\left(\frac{g^{2}}{2}-f^{2}_{1}\right)(\eta^{\dagger}\rho)(\rho^{\dagger}\eta)+\left(\frac{g^{2}}{2}-f^{2}_{1}\right)(\eta^{\dagger}\chi)(\chi^{\dagger}\eta)+\left(\frac{g^{2}}{2}-f^{2}_{1}\right)(\rho^{\dagger}\chi)(\chi^{\dagger}\rho)$
-- $\frac{g^{2}}{6}\left[(\eta^{\dagger}\eta)Tr[(S^{\dagger}S)]+(\rho^{\dagger}\rho)Tr[(S^{\dagger}S)]+(\chi^{\dagger}\chi)Tr[(S^{\dagger}S)]\right]$
$+$ $\frac{g^{2}}{2}Tr[(S^{\dagger}\eta)(\eta^{\dagger}S)]+\left(\frac{g^{2}}{2}+f^{2}_{3}\right)\left\{Tr[(S^{\dagger}\rho)(\rho^{\dagger}S)]+Tr[(S^{\dagger}\chi)(\chi^{\dagger}S)]\right\}$
$+$ $\left(\mu^{2}_{S}-\frac{g^{2}}{2}\right)\left(Tr[(S^{\dagger}S)]\right)^{2}+g^{2}Tr[(S^{\dagger}S)(S^{\dagger}S)]$
$+$ $\left\{f_{1}f_{3}\left[\epsilon(\rho^{\dagger}\eta^{\dagger})(\rho S)+\epsilon(\chi^{\dagger}\eta^{\dagger})(\chi S)\right]+H.c.\right\}+\ldots$
and comparing it with Eq.(200) we can see
$\mu^{2}_{1}$ $=$ $m^{2}_{\eta},\quad\mu^{2}_{2}=m^{2}_{\rho},\quad\mu^{2}_{3}=m^{2}_{\chi},\quad\mu^{2}_{4}=m^{2}_{S},$
$f_{1}$ $=$ $-k_{1},\quad f_{2}=-k_{3},$
$a_{1}$ $=$ $\mu^{2}_{\eta}-\frac{g^{2}}{6},\quad e_{1}=\mu^{2}_{S}-\frac{g^{2}}{2},\quad e_{2}=g^{2},$
$a_{2}$ $=$ $\mu^{2}_{\rho}+\frac{g^{\prime 2}}{2}-\frac{g^{2}}{6},a_{3}=\mu^{2}_{\chi}+\frac{g^{\prime 2}}{2}-\frac{g^{2}}{6},$
$a_{4}$ $=$ $-\frac{g^{2}}{6}+\frac{f^{2}_{1}}{9}=a_{6}=a_{7}=a_{8}=a_{9},$
$a_{5}$ $=$ $-g^{\prime 2}-\frac{g^{2}}{6}+\frac{f^{2}_{1}}{9},$
$d_{1}$ $=$ $-\frac{g^{2}}{6}=d_{3}=d_{4}=d_{5},$
$d_{2}$ $=$ $d_{4}=d_{6}=\frac{g^{2}}{2},$
$a_{10}$ $=$ $=0.$ (217)